### v2

# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part II: Information Theoretic Constraints Breed New Combinatorial Structures

Abstract
The characterization of entropy functions is of fundamental importance in information theory. By imposing constraints on their Shannon outer bound, i.e., the polymatroidal region, one obtains the faces of the region and entropy functions on them with special structures. In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of the polymatroidal region $\Gamma_{4}$. In Part I, we formulated the problem, enumerated all $59$ types of $2$-dimensional faces of $\Gamma_{4}$ by an algorithm, and fully characterized entropy functions on $49$ types of them. In this paper, i.e., Part II, we will characterize entropy functions on the remaining $10$ types of faces, among which $8$ types are fully characterized, and $2$ types are partially characterized. To characterize these types of faces, we introduce some new combinatorial design structures that are interesting in themselves.
Index terms-entropy function, polymatroid, information inequalities, orthogonal array, multi-level variable-strength orthogonal array
I Introduction
Let $N_{n}=\{1,2,\ldots,n\}$ and $\mathbf{X}\triangleq(X_{i},i\in N_{n})$ be a random vector indexed by $N_{n}$. The set function $\mathbf{h}:2^{N_{n}}\to\mathbb{R}$ defined by
$\mathbf{h}(A)=H(X_{A}),\quad A\subseteq N_{n}$
is called the entropy function of $\mathbf{X}$, while $\mathbf{X}$ is called a characterizing random vector of $\mathbf{h}$. The Euclidean space $\mathcal{H}_{n}\triangleq\mathbb{R}^{2^{N_{n}}}$ where entropy functions live is called the entropy space of degree $n$. The set of all entropy functions, denoted by $\Gamma^{*}_{n}$, is called the entropy region of degree $n$. The characterization of entropy functions, i.e., determining whether an $\mathbf{h}\in\mathcal{H}_{n}$ is in $\Gamma^{*}_{n}$, is of fundamental importance in information theory.
Entropy functions are (the rank functions of) polymatroids, i.e., they satisfy polymatroidal axioms, that is, for all $A,B\subseteq N_{n}$,
$\mathbf{h}(A)$ $\geq 0,$ (I.1)
$\mathbf{h}(A)$ $\leq\mathbf{h}(B),\quad\text{ if }A\subseteq B,$ (I.2)
$\mathbf{h}(A)+\mathbf{h}(B)$ $\geq\mathbf{h}(A\cap B)+\mathbf{h}(A\cup B).$ (I.3)
The region in $\mathcal{H}_{n}$ bounded by such inequalities, denoted by $\Gamma_{n}$ is called the polymatroidal region of degree $n$. Thus, $\Gamma_{n}$ is an outer bound on $\Gamma^{*}_{n}$. For more about entropy functions, we referred the readers to [1, Chapter 13-15].
Traditionally, entropy functions are characterized by information inequalities. Those inequalities derived by polymatroidal axioms are called Shannon-type, as they correspond to the non-negativity of Shannon information measures. Since 1998, a series of non-Shannon-type information inequalities, among which Zhang-Yeung inequality is the first one [2], were discovered [3][4][5]. Thus $\overline{\Gamma^{*}_{n}}$, the closure of $\Gamma^{*}_{n}$, is strictly included in $\Gamma_{n}$ when $n\geq 4$. In 2007, Matúš showed that for $n\geq 4$, there exist infinitely non-redundant non-Shannon-type inequalities, indicating that $\overline{\Gamma_{n}^{*}}$ is not polyhedral [6]. Each information inequality determines an outer bound on $\Gamma^{*}_{n}$, as those set functions in $\mathcal{H}_{n}$ that dissatisfy it must be located outside $\Gamma^{*}_{n}$. In this series of two papers, we develop a system of entropy function characterization from the perspective of faces of $\Gamma_{n}$, which covers the traditional inequality characterization.
By definition, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$. Thus, each Shannon-type information inequality determines a face $F$ of $\Gamma_{n}$. It is natural to characterize entropy functions on the specific $F$ of $\Gamma_{n}$ (See Subsection II-B for details on the faces of a polyhedral cone). Let $F^{*}\triangleq F\cap\Gamma^{*}_{n}$ be the set of all entropy functions in $F$. In the following, to determine the entropy functions on $F$, or the region $F^{*}$, we will call it characterize $F$ for short. A non-Shannon-type information inequality can be considered as an outer bound on $F^{*}$ when $F=\Gamma_{n}$ itself, the improper face of $\Gamma_{n}$. A constrained non-Shannon-type information inequality is an outer bound on $F^{*}$ when $F$ is the face determined by the constraints that are equalities obtained by setting some Shannon-type inequalities as equalities. When $F$ is an extreme ray, i.e., a $1$-dimensional face of $\Gamma_{n}$, if it contains a matroid, entropy functions on $F$ are called matroidal entropy functions, and they can be fully characterized by the probabilistically characteristic set of the matroid [7][8][9]. Matúš fully characterized the first non-trivial 2-dimensional face of $\Gamma_{n}$ in 2006 [10]. It is a $2$-dimensional face of $\Gamma_{3}$. In 2012, Chen and Yeung characterized another $2$-dimensional face of $\Gamma_{3}$ [11]. They are the only two types of non-trivial faces of $\Gamma_{3}$ that need to be characterized. To the best of our knowledge, so far, there is no fully characterized no-trivial $3$-dimensional faces. However, partial characterizations of $3$-dimensional faces of $\Gamma_{3}$ can be found in [12][13].
Many information-theoretic problems can be considered as applications of entropy function characterizations on faces of $\Gamma_{n}$. In a series of three papers [14, 15, 16], Matúš and Studneý solved the probabilistic conditional independence problem for four random variables. Note each class of conditional independence constraints, which is called a semimatroid in their papers, determines a face $F$ of $\Gamma_{n}$. The solution to this problem, i.e., the probabilistic-representability of a semimatroid determines whether the relative interior of the corresponding face $F$ intersects with $\Gamma^{*}_{n}$. This problem thus can be considered as partial characterizations of the faces of $\Gamma_{n}$. In [17], Yan, Yeung, and Zhang proved a formula involving $\Gamma_{n}^{*}$ for the capacity of multi-source multi-sink network coding. Those constraints in the formula induced by network topology and source independency form a face of $\Gamma_{n}$, which shows that this holy-grail network coding problem corresponds to the entropy function characterizations on this face. For the secret-sharing problem, see [18] for example, the perfect secrecy and decoding correctness conditions of an access structure determine a face $F$ on $\Gamma_{n}$, and the information ratio of the secret-sharing problem can be considered as an optimization problem whose feasible region is $F\cap\Gamma^{*}_{n}$. Other problems, such as distributed data storage [19], coded caching [20], Markov random fields [21], and relational database [22] are also related to the entropy regions on the faces of $\Gamma_{n}$.
Though the information theory problems discussed above usually involve more than four random variables, and the corresponding faces are of dimension higher than $2$, following the characterizations of extreme rays. i.e., $1$-dimensional faces containing matroid [7][8][9], and $2$-dimensional faces of $\Gamma_{3}$ [10][11], in this series of two papers, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$, which may serve as stepping stones to the general cases of this problem. In Part I [23], we enumerated all $59$ types $2$-dimensional faces of $\Gamma_{4}$ by an algorithm and completely characterized $49$ types of them. In this part, we characterize the remaining $10$ types of faces, among which $8$ types are fully characterized and $2$ types are partially characterized. To characterize these types of faces, we adopt two sets of combinatorial structures, that is, mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes, and the characterizations breed some new combinatorial design structures that are interesting in themselves.
The rest of this paper is organized as follows. In Section II, for self-contained, we give the preliminaries on integer polymatroids and matroids, polyhedral cones, mixed-level variable-strength orthogonal arrays and their relationship to orthogonal Latin hypercubes. We then list the results of the characterization of extreme rays of $\Gamma_{4}$, which has been done in Part I, without proofs. In Section III, we complete the characterization of the remaining $10$ types of $2$-dimensional faces of $\Gamma_{4}$, and the results are summarized in Table LABEL:table:2 together with those in Part I. The correspondences between the faces and the theorems characterizing them in two papers are summarized in Table 4.
II Preliminaries
II-A Integer polymatroids and matroids
For a polymatroid $\mathbf{h}\in\Gamma_{n}$, if $\mathbf{h}(A)\in\mathbb{Z}$ for all $A\subseteq N_{n}$, $\mathbf{h}$ is called integer. An ordered pair $M=(N_{n},\mathbf{r})$ is called a matroid with rank function $\mathbf{r}$ if $\mathbf{r}$ is an integer polymatroid with $\mathbf{r}(A)\leq|A|$ for all $A\subseteq N_{n}$. Like polymatroids, in this paper, we do not distinguish a matroid and its rank function unless otherwise specified.
A uniform matroid $U_{k,n}$ is a matroid with rank function $\mathbf{r}(A)=\min\{k,|A|\}$ for all $A\subseteq N_{n}$.
For a matroid $M$ and $e\in N_{n}$, if $\mathbf{r}(e)=0$, $e$ is called a loop of $M$. For $e,e^{\prime}\in N_{n}$, if $\mathbf{r}(\{e,e^{\prime}\})=1$, then $e$ and $e^{\prime}$ are called parallel.
For more about matroid theory, readers are referred to [24].
II-B Faces of a polyhedral cone
Let $C\subseteq\mathbb{R}^{d}$ be a full-dimensional polyhedral cone. For a hyperplane $P$ containing the origin $O$ in $\mathbb{R}^{d}$, if $C\subseteq P^{+}$, where $P^{+}$ is one of the two halfspaces corresponding to $P$, $F\triangleq C\cap P$ is called a (proper) face of $C$, while $C$ itself is its improper face. When $\dim F=d-1$, $F$ is called a facet of $C$, and when $\dim F=1$, $F$ is an extreme ray of $C$. Either the set of all facets or the set of all extreme rays of $C$ uniquely determines the cone, and they are called the H-representation and $V$-representation of the cone, respectively. For each face $F$ of the cone, it can be written as the intersection of the facets of the cone that contain $F$, or the convex hull of the extreme rays contained in $F$. We also call them the H-representation and V-representation of $F$, respectively. More details on this topic are referred to [25].
As we discussed in Section I, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$ determined by polymatroidal axioms in (I.1)-(I.3). They are equivalent to the following elemental inequalities
$\mathbf{h}(N_{n})\geq\mathbf{h}({N_{n}\setminus\{i\}})\quad\quad i\in N_{n};$ (II.2.1)
$\mathbf{h}(K)+\mathbf{h}(K\cup ij)\leq\mathbf{h}(K\cup i)+\mathbf{h}(K\cup j),$ (II.2.2)
$\quad\quad i,j\in N_{n},K\subseteq N_{n}\setminus\{i,j\}$
each of which determines a facet of $\Gamma_{n}$[1, Chapter 14]. When $n=4$, it can be checked that there are $28$ elemental inequalities, or $28$ facets of $\Gamma_{4}$.
It can be seen in [26] that there are $41$ extreme rays of $\Gamma_{4}$. Note that each extreme ray $E$ of $\Gamma_{4}$ can be written in the form
$E=\{a\mathbf{r}:a\geq 0\}$ (II.2.3)
where $\mathbf{r}$ is the minimal integer polymatroid in the ray , that is, an integer polymatroid such that $\mathbf{r}/t$ is not integer for any integer $t>1$. Therefore, in this paper, when we say a minimal integer polymatroid, we mean the extreme ray containing it unless otherwise specified. Note that some of these integer polymatroids are matroids. The $41$ extreme rays can be classified into the following $11$ types.
  • $U^{i}_{1,1}$, $i\in N_{4}$;
  • $U^{\alpha}_{1,2}$, $\alpha\subseteq N_{4}$, $|\alpha|=2$;
  • $U^{\alpha}_{1,3}$, $\alpha\subseteq N_{4}$, $|\alpha|=3$;
  • $U^{\alpha}_{2,3}$, $\alpha\subseteq N_{4}$, $|\alpha|=3$;
for $U^{\alpha}_{k,m}$ with $\alpha\subseteq N_{4}$ and $|\alpha|=m$, we mean a matroid on $N_{4}$ whose restriction on $\alpha$ is a $U_{k,m}$ and $e\in N_{4}\setminus\alpha$ being loops;
  • $U_{1,4}$;
  • $\mathcal{W}^{\alpha}_{2}$, $\alpha\subseteq N_{4}$, $|\alpha|=2$;
for $\mathcal{W}^{\alpha}_{2}$ with $\alpha\subseteq N_{4}$, $|\alpha|=2$, it is called a wheel matroid with order 22, [^1] We adopt the notation and terminology in [24, Section 8.4] and it is a matroid with two parallel elements in $\alpha$, and each element in $\alpha$ and the other two elements in $N_{4}$ form a $U_{2,3}$;
  • $U_{2,4}$;
  • $U_{3,4}$;
  • $\hat{U}^{i}_{2,5}$, $i\in N_{4}$;
for $\hat{U}^{i}_{2,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{2,5}$, [^2] See [27, Theorem 1.3.6] and [28, Theorem 4] for the definition of free expansion. and its rank function is defined by
$\mathbf{r}(A)=\begin{cases}\min\{2,|A|\},\quad&A\neq\{i\}\\ 2,\quad&A=\{i\}\end{cases}$
for any $A\subseteq N_{4}$,
  • $\hat{U}^{i}_{3,5}$;
for $\hat{U}^{i}_{3,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{3,5}$, and its rank function is defined by
$\mathbf{r}(A)=\begin{cases}\min\{3,|A|+1\},\quad&i\in A,\\ |A|,\quad&i\not\in A,\end{cases}$
for any $A\subseteq N_{4}$;
  • $V^{\alpha}_{8}$, $\alpha\subseteq N_{4}$, $|\alpha|=2$;
for $V^{\alpha}_{8}$ with $\alpha\subseteq N_{4}$ and $|\alpha|=2$, it is a polymatroid whose free expansion is the Vámos matroid, and its rank function is defined by
$\mathbf{r}(A)=\begin{cases}3,\quad&|A|=2\text{ and }A\neq\alpha,\\ \min\{4,2|A|\},\quad&\text{ o.w.}\end{cases}$
It can be seen that for an extreme ray in the form $E^{\alpha}$ with $\alpha\subseteq N_{4}$, it is in a type with $\binom{4}{|\alpha|}$ extreme rays, and each extreme ray in the type can be obtained from each other by permuting the indices in $N_{4}$. For a specific extreme ray, say $U^{\{1,2\}}_{1,2}$, for simplicity, we will drop the bracket and comma of the set in the superscript, and write it as $U^{12}_{1,2}$. To facilitate the readers, the $11$ types of extreme rays are presented in Table 1 by the rank functions of their minimal integer polymatroid in the form of $15$-dimensional vectors, i.e., $\mathbf{r}=(\mathbf{r}(A),\emptyset\neq A\subseteq N_{4})$.
In the same manner, we denote the $8$ extreme rays of $\Gamma_{3}$ and classify them in the following $4$ types.
  • $U^{i;3}_{1,1}$, $i\in N_{3}$;
  • $U^{\alpha;3}_{1,2}$, $\alpha\subseteq N_{3}$, $|\alpha|=2$;
  • $U_{1,3}$;
  • $U_{2,3}$.
Here we put second superscript $3$ to $U^{i;3}_{1,1}$ and $U^{\alpha;3}_{1,1}$ to distinguish them from $U^{i}_{1,1}$ and $U^{\alpha}_{1,1}$, the extreme rays of $\Gamma_{4}$, respectively.
Table 1: Extreme rays of $\Gamma_{4}$ and their rank functions
[…]
$U^{1}_{1,1}$ 11 0 0 0 11 11 11 0 0 0 11 11 11 11 11
$U^{12}_{1,2}$ 11 11 0 0 11 11 11 11 11 0 11 11 11 11 11
$U^{123}_{1,3}$ 11 11 11 0 11 11 11 11 11 11 11 11 11 11 11
$U_{1,4}$ 11 11 11 11 11 11 11 11 11 11 11 11 11 11 11
$U^{123}_{2,3}$ 11 11 11 0 22 22 11 22 11 11 22 22 22 22 22
$\mathcal{W}^{12}_{2}$ 11 11 11 11 11 22 22 22 22 22 22 22 22 22 22
$U_{2,4}$ 11 11 11 11 22 22 22 22 22 22 22 22 22 22 22
$U_{3,4}$ 11 11 11 11 22 22 22 22 22 22 33 33 33 33 33
$\hat{U}^{1}_{2,5}$ 22 11 11 11 22 22 22 22 22 22 22 22 22 22 22
$\hat{U}^{1}_{3,5}$ 22 11 11 11 33 33 33 22 22 22 33 33 33 33 33
$V^{12}_{8}$ 22 22 22 22 44 33 33 33 33 33 44 44 44 44 44
II-C Mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes
To characterize entropy functions on the extreme rays of $\Gamma_{4}$, we adopt two sets of combinatorial structures, that is, mixed-level variable-strength orthogonal arrays and orthogonal Latin hypercubes, which are equivalent in some sense, but each has its own advantages. For mixed-level variable-strength orthogonal arrays, it will be generalized more easily to the cases with more random variables and higher dimensions, and the symmetries between random variables are more straightforward. In contrast, the language of orthogonal Latin hypercubes will be better visualized for the cases of four random variables. Furthermore, they will also play important roles in characterizing $2$-dimensional faces of $\Gamma_{4}$, especially those with two extreme rays containing integer polymatroids of rank exceeding $1$.
Definition 1 (MVOA, [9]).
Given an integer polymatroid $P=(N_{n},\mathbf{r})$ with $\mathbf{r}(N_{n})\geq 2$ and an integer $v>1$, a $v^{\mathbf{r}(N_{n})}\times n$ array $\mathbf{T}$ with columns indexed by $N_{n}$, whose entries of a column indexed by $i\in N_{n}$ are from a set $I_{i}$ with cardinality $v^{\mathbf{r}(i)}-1$, is called a mixed-level variable-strength orthogonal array(MVOA) induced by $P$ with base level $v$ if for each $A\subseteq N_{n}$, $v^{\mathbf{r}(N_{n})}\times|A|$ subarray $\mathbf{T}(A)$ of $\mathbf{T}$ consisting of columns indexed by $A$ satisfies the following condition: each row of $\mathbf{T}(A)$ occurs in $\mathbf{T}(A)$ exactly $v^{\mathbf{r}(N_{n})-\mathbf{r}(A)}$ times. We also call such $\mathbf{T}$ a $\mathrm{MVOA}(P,v)$.
Usually, we set $I_{i}=\mathbb{I}_{v^{\mathbf{r}(i)}-1}$, unless otherwise specified, where $\mathbb{I}_{u}\triangleq\{0,1,\ldots,u-1\}$ for any integer $u\geq 1$. When the integer polymatroid is a matroid $M$, its induced MVOA is reduced to a variable-strength orthogonal array $\mathrm{VOA}(M,v)$[7],[8],[9]. When a matroid is a uniform matroid $U_{t,n}$, the induced $\mathrm{VOA}(U_{t,n},v)$ is an orthogonal array with index unity, usually denoted by $\mathrm{OA}(t,n,v)$ [7, 8].
To describe MVOAs, we introduce some notations which will be used throughout this paper. The $v^{\mathbf{r}(N_{n})}\times|A|$ subarray of $\mathbf{T}$ consisting of columns indexed by $A$ is denoted by $\mathbf{T}(A)$. For simplicity, we drop the brackets of $\mathbf{T}(A)$, e.g., we write $\mathbf{T}(\{1,2,3\})$ as $\mathbf{T}(1,2,3)$. Let $\mathbf{T}(A;j)$ denote the $j$-th row of $\mathbf{T}(A)$.
For the orthogonal Latin hypercubes, we will not give a general definition in this paper. Because, on the one hand, for some extreme rays containing an integer polymatroid other than a uniform matroid, the variants of traditional orthogonal Latin hypercubes are utilized. On the other hand, in this paper, we only consider the cases corresponding to integer polymatroids in $\Gamma_{4}$, and so, a case-by-case discussion is enough. For the general form of these variants, it can be studied in future research.
We begin our discussion from Latin squares, the non-trivial orthogonal Latin hypercubes with the smallest parameters, which corresponds to the smallest connected matroid with rank exceeding $1$. For uniform matroid $U_{2,3}$, a $\mathrm{VOA}(U_{2,3},v)$ corresponds to a Latin square $S$ of order $v$ with row and column indices and symbols in $\mathbb{I}_{v}$, that is, a $v\times v$ square $S$ with cells indexed by $(i,j)\in\mathbb{I}^{2}_{v}$ and symbols in the cells of each row $\{S(i,j),j\in\mathbb{I}_{v}\}$ and each column $\{S(i,j),i\in\mathbb{I}_{v}\}$ all distinct. For each row $(i,j,k)$ of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$, it corresponds to the cell $S(i,j)$ with symbol $k$. Here, we also define Latin squares of the zeroth-kind and the second-kind, [^3] The definition of Latin squares of the zeroth-kind and the second-kind are in the spirit of the definition of the “Latin cubes of first order and second order” in [29]. “Order” is usually used for the size of a Latin square, so in our paper, we use “kind” instead. In the following, Latin cubes of the first kind and the second kind will be defined similarly. which are trivial themselves, but can be considered as building blocks in construction of other combinatorial structures in this paper, while for a traditional Latin square, we call it of the first kind. For a $v\times v$ square $S$ with symbols in all cells, each identical to 0, we call it a Latin square of the zeroth kind. It corresponds to a $\mathrm{VOA}(U^{1,2;3}_{1,1},v)$, where $U_{1,1}^{1,2;3}$ is a matroid whose rank function is the summation of the rank functions of $U^{1;3}_{1,1}$ and $U^{2;3}_{1,1}$. For a $v\times v$ square $S$ with symbols in all cells distinct from $\mathbb{I}_{v^{2}}$, we call it a Latin square of the second kind. It corresponds to a $\mathrm{MVOA}(U^{13,23;3}_{1,2},v)$, where $U^{13,23;3}_{1,2}$ is an integer polymatroid whose rank function is the summation of the rank functions of $U^{13;3}_{1,2}$ and $U^{23;3}_{1,2}$. For a traditional Latin square, or a Latin square of the first kind, we call it a Latin square and omit “of the first kind” if there is no ambiguity.
Example 1.
In Table 2, we list Latin squares of the zeroth kind, the first kind, and the second kind, and their corresponding MVOAs for $v=3$.
Table 2: Three kinds of Latin squares and their corresponding (M)VOAs
(a) (a) The zeroth kind
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&0&0&0\\ 1&0&0&0\\ 2&0&0&0\end{array}$
$\mathrm{VOA}(U^{1,2;3}_{1,1},3):$ $\begin{array}[]{ccc}0&0&0\\ 0&1&0\\ 0&2&0\\ 1&0&0\\ 1&1&0\\ 1&2&0\\ 2&0&0\\ 2&1&0\\ 2&2&0\end{array}$
(b) (b) The first kind
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&0&1&2\\ 1&2&0&1\\ 2&1&2&0\end{array}$
$\mathrm{VOA}(U_{2,3},3):$ $\begin{array}[]{ccc}0&0&0\\ 0&1&1\\ 0&2&2\\ 1&0&2\\ 1&1&0\\ 1&2&1\\ 2&0&1\\ 2&1&2\\ 2&2&0\end{array}$
(c) (c) The second kind
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&0&1&2\\ 1&3&4&5\\ 2&6&7&8\end{array}$
$\mathrm{MVOA}(U^{13,23;3}_{1,2},3):$ $\begin{array}[]{ccc}0&0&0\\ 0&1&1\\ 0&2&2\\ 1&0&3\\ 1&1&4\\ 1&2&5\\ 2&0&6\\ 2&1&7\\ 2&2&8\end{array}$
In the following, we list the mixed-level variable-strength orthogonal array induced by integer polymatroids on the extreme rays of $\Gamma_{4}$ with rank exceeding $1$, and their corresponding variants of orthogonal Latin hypercubes. For each type of extreme rays, we only consider one of its representatives.
  • For a $\mathrm{VOA}(U^{123}_{2,3},v)$ $\mathbf{T}$, $\mathbf{T}(1,2,3)$ forms a $\mathrm{VOA}(U_{2,3},v)$, and all the entries of the column $\mathbf{T}(4)$ can be a constant symbol 0. It corresponds to a pair of Latin squares sharing the same row and column indices and one of which is the first-kind, and the other of which is the zeroth-kind, where $\mathbf{T}(1)$ and $\mathbf{T}(2)$, $\mathbf{T}(3)$, $\mathbf{T}(4)$ corresponds to the indices of rows and columns, the symbols of the first square and symbols of the second square, respectively.
  • For a $\mathrm{VOA}(\mathcal{W}^{34},v)$ $\mathbf{T}$, $\mathbf{T}((1,2,3)$ forms a $\mathrm{VOA}(U_{2,3},k)$ and $\mathbf{T}(4)=\mathbf{T}(3)$. Thus $\mathbf{T}$ corresponds to a pair of $v\times v$ identical Latin squares.
  • For a $\mathrm{VOA}(U_{2,4},v)$ $\mathbf{T}$, it correspond to a pair of mutual orthogonal $v\times v$ Latin squares. A pair of Latin squares (of the first kind) $S_{1}$ and $S_{2}$ are called orthogonal if the pairs of symbols in the pairs of cells $\{(S_{1}(i,j),S_{2}(i,j)):i,j\in\mathbb{I}_{v}\}$ are all distinct, or exactly those in $\mathbb{I}^{2}_{v}$.
  • For a $\mathrm{VOA}(U_{3,4},v)$ $\mathbf{T}$, it corresponds to a Latin cube of order $v$ (type 2 and the first kind, permutation cube [30, VI.22.33]), where entries in $\mathbf{T}(1)$, $\mathbf{T}(2)$ and $\mathbf{T}(3)$ correspond to the indices of three dimensions, respectively, and $\mathbf{T}(4)$ corresponds to the symbols. A Latin cube of order $v$ and the first kind is a $v\times v\times v$ cube with symbols in each line $\{C(i,j,k),i\in\mathbb{I}_{v}\}$ for fixed $j,k\in\mathbb{I}_{v}$, $\{C(i,j,k),j\in\mathbb{I}_{v}\}$ for fixed $i,k\in\mathbb{I}_{v}$, and $\{C(i,j,k),i\in\mathbb{I}_{v}\}$ for fixed $i,j\in\mathbb{I}_{v}$ all distinct.
  • For a $\mathrm{MVOA}(\hat{U}^{4}_{2,5},v)$ $\mathbf{T}$, it corresponds to a pair of $v\times v$ squares $(S_{1},S_{2})$, where $S_{1}$ is a Latin square of the first kind, and $S_{2}$ is a Latin square of the second kind.
  • For a $\mathrm{MVOA}(\hat{U}^{4}_{3,5},v)$ $\mathbf{T}$, it corresponds a Latin cube $C$ of order $v$ and the second kind. A Latin cube of order $v$ and the second kind is $v\times v\times v$ cube with each layer $\{C(i,j,k),i,j\in\mathbb{I}_{v}\}$ for fixed $k\in\mathbb{I}_{v}$, $\{C(i,j,k),i,k\in\mathbb{I}_{v}\}$ for fixed $j\in\mathbb{I}_{v}$, and $\{C(i,j,k),j,k\in\mathbb{I}_{v}\}$ for fixed $i\in\mathbb{I}_{v}$ all Latin squares of order $v$ and the second kind.
It can be seen that MVOAs induced by rank $2$ integer ploymatroids correspond to a pair of Latin squares of the same kind superimposed in different ways, or of different kinds. While MVOAs induced by $2$ rank $3$ integer ploymatroids correspond to Latin cubes of two different kinds.
Example 2.
In Table 3, we list orthogonal Latin hypercubes and their corresponding MVOAs discussed above for $v=3$.
Table 3: Orthogonal Latin hypercubes and their corresponding MVOAs
(a) (a) Latin squares of the first and zeroth kinds
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&(0,0)&(1,0)&(2,0)\\ 1&(2,0)&(0,0)&(1,0)\\ 2&(1,0)&(2,0)&(0,0)\end{array}$
$\mathrm{VOA}(U^{123}_{2,3},3):$ $\begin{array}[]{cccc}0&0&0&0\\ 0&1&1&0\\ 0&2&2&0\\ 1&0&2&0\\ 1&1&0&0\\ 1&2&1&0\\ 2&0&1&0\\ 2&1&2&0\\ 2&2&0&0\end{array}$
(b) (b) Two orthogonal Latin squares
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&(0,0)&(1,1)&(2,2)\\ 1&(2,1)&(0,2)&(1,0)\\ 2&(1,2)&(2,0)&(0,1)\end{array}$
$\mathrm{VOA}(U_{2,4},3):$ $\begin{array}[]{cccc}0&0&0&0\\ 0&1&1&1\\ 0&2&2&2\\ 1&0&2&1\\ 1&1&0&2\\ 1&2&1&0\\ 2&0&1&2\\ 2&1&2&0\\ 2&2&0&1\end{array}$
(c) (c) Two identical Latin squares
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&(0,0)&(1,1)&(2,2)\\ 1&(2,2)&(0,0)&(1,1)\\ 2&(1,1)&(2,2)&(0,0)\end{array}$
$\mathrm{VOA}(\mathcal{W}^{34}_{2},3):$ $\begin{array}[]{cccc}0&0&0&0\\ 0&1&1&1\\ 0&2&2&2\\ 1&0&2&2\\ 1&1&0&0\\ 1&2&1&1\\ 2&0&1&1\\ 2&1&2&2\\ 2&2&0&0\end{array}$
(d) (d) Latin squares of the first and second kinds
$\begin{array}[]{c|ccc}&0&1&2\\ \hline\cr 0&(0,0)&(1,1)&(2,2)\\ 1&(2,3)&(0,4)&(1,5)\\ 2&(1,6)&(2,7)&(0,8)\end{array}$
$\mathrm{MVOA}(\hat{U}^{4}_{2,5},3):$ $\begin{array}[]{cccc}0&0&0&0\\ 0&1&1&1\\ 0&2&2&2\\ 1&0&2&3\\ 1&1&0&4\\ 1&2&1&5\\ 2&0&1&6\\ 2&1&2&7\\ 2&2&0&8\end{array}$
(e) (e) Latin cube of the first kind
$\begin{array}[]{c|ccc|ccc|ccc}&0&1&2&0&1&2&0&1&2\\ \hline\cr 0&(0,0)&(0,1)&(0,2)&(1,1)&(1,2)&(1,0)&(2,2)&(2,0)&(2,1)\\ 1&(0,1)&(0,2)&(0,0)&(1,2)&(1,0)&(1,1)&(2,0)&(2,1)&(2,2)\\ 2&(0,2)&(0,0)&(0,1)&(1,0)&(1,1)&(1,2)&(2,1)&(2,2)&(2,0)\end{array}$
$\mathrm{VOA}(U_{3,4},3)$ (transposed):
[…]
(f) (f) Latin cube of the second kind
$\begin{array}[]{c|ccc|ccc|ccc}&0&1&2&0&1&2&0&1&2\\ \hline\cr 0&(0,0)&(0,1)&(0,2)&(1,4)&(1,5)&(1,6)&(2,8)&(2,3)&(2,7)\\ 1&(0,3)&(0,4)&(0,5)&(1,7)&(1,8)&(1,0)&(2,2)&(2,6)&(2,1)\\ 2&(0,6)&(0,7)&(0,8)&(1,1)&(1,2)&(1,3)&(2,5)&(2,0)&(2,4)\end{array}$
$\mathrm{MVOA}(\hat{U}^{4}_{3,5},3)$ (transposed):
[…]
II-D Entropy functions on the extreme rays of $\Gamma_{4}$
As we discussed in Subsection II-B, for the V-representation of a face, it can be written as the convex hull of the extreme rays it contains. Therefore, to characterize entropy functions on $2$-dimensional faces of $\Gamma_{4}$, we need first to characterize the entropy functions on the extreme rays of $\Gamma_{4}$. For the eleven types of extreme rays, the entropy functions on seven of them are matroidal and have been characterized in [7], while those containing $\hat{U}^{i}_{2,5}$ or $\hat{U}^{i}_{3,5}$ are characterized in Part I of this series [23], and those containing $V^{\alpha}_{8}$ are non-entropic. Theorems that characterize these entropy functions are listed below without proof.
Theorem II-D.1.
For $E=U^{1}_{1,1},U^{12}_{1,2},U^{123}_{1,3},U_{1,4}$, i.e., extreme rays containing a matroid with rank $1$, $a\mathbf{r}\in E$ is entropic for all $a\geq 0$.
The characterization of the following four types of extreme rays follows immediately from matroidal entropy functions in [7] and [9].
Theorem II-D.2.
For $E=U^{123}_{2,3},\mathcal{W}^{14}_{2},U_{3,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log k$ for integer $k\geq 1$.
Theorem II-D.3.
([7, Proposition 2]) For $E=U_{2,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log k$ for positive integer $k\neq 2,6$.
Theorem II-D.4.
([23, Theorem II-C.4]) For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
Theorem II-D.5.
([23, Theorem II-C.5]) For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log k$ for integer $k>0$.
Theorem II-D.6.
For $E=V^{\alpha}$, $a\mathbf{h}\in E$ is entropic if and only if $a=0$.
Remark In Part I [23], Theorem II-D.4(resp. II-D.5) on extreme ray containing $\hat{U}^{i}_{2,5}$(resp. $\hat{U}^{i}_{3,5}$) was proved by the construction of a specific $\mathrm{MVOA}(\hat{U}^{i}_{2,5},v)$(resp. $\mathrm{MVOA}(\hat{U}^{i}_{3,5},v)$) for all $v\geq 1$. However, the correspondence between a pair of Latin squares of the first kind and the second kind and an $\mathrm{MVOA}(\hat{U}^{i}_{2,5},v)$(resp. a Latin cube of the second kind and an $\mathrm{MVOA}(\hat{U}^{i}_{3,5},v)$) setting up in Subsection II-C provides a general construction of the problem. Such construction also sheds some light on the open question [9, Question 1].
Table 4: Two-dimensional faces of $\Gamma_{4}$.
$U^{i}_{1,1}$ $U^{\alpha}_{1,2}$ $U^{\alpha}_{1,3}$ $U_{1,4}$ $U^{\alpha}_{2,3}$ $\mathcal{W}^{\alpha}_{2}$ $U_{2,4}$ $U_{3,4}$ $\hat{U}^{i}_{2,5}$ $\hat{U}^{i}_{3,5}$ $V^{\alpha}_{8}$
$U^{j}_{1,1}$ $(U^{12}_{1,2},U^{1}_{1,1}),12$ $(U^{123}_{1,3},U^{1}_{1,1}),12$ $(U_{2,3}^{123},U_{1,1}^{1}),12$ $(\mathcal{W}_{2}^{14},U_{1,1}^{1}),12$ $(\hat{U}^{1}_{2,5},U_{1,1}^{1}),4$ $(\hat{U}^{1}_{3,5},U^{1}_{1,1}),4$ $(V^{12}_{8},U^{1}_{1,1}),12$
$(U^{1}_{1,1},U^{2}_{1,1}),6$ Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk1 $(U_{1,4},U^{1}_{1,1}),4$ Part 1 Thm.LABEL:part1-rk3 Part 1 Thm.LABEL:part1-rk3 $(U_{2,4},U^{1}_{1,1}),4$ $(U_{3,4},U^{1}_{1,1}),4$ Part 1 Thm.LABEL:part1-rk18 Part 1 Thm.LABEL:part1-rk21 Part 1 Thm.LABEL:part1-rk31
Part 1 Thm.LABEL:part1-rk1 $(U^{12}_{1,2},U^{3}_{1,1}),12$ $(U^{123}_{1,3},U^{4}_{1,1}),4$ Part 1 Thm.LABEL:part1-rk1 $(U_{2,3}^{123},U_{1,1}^{4}),4$ $(\mathcal{W}_{2}^{34},U_{1,1}^{1}),12$ Part 2 Thm.III-A.3 Part 1 Thm.LABEL:part1-rk8 $(\hat{U}^{1}_{2,5},U_{1,1}^{2}),12$ $(\hat{U}^{1}_{3,5},U^{2}_{1,1}),12$ $(V^{12}_{8},U^{3}_{1,1}),12$
Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk3 Part 1 Thm.LABEL:part1-rk3 Part 1 Thm.LABEL:part1-rk18 Part 1 Thm.LABEL:part1-rk21 Part 1 Thm.LABEL:part1-rk31
$U^{\beta}_{1,2}$ $\backslash$ $(\mathcal{W}_{2}^{14},U_{1,2}^{14}),6$
$(U^{12}_{1,2},U^{13}_{1,2}),12$ $(U^{123}_{1,3},U^{12}_{1,2}),12$ $(U_{2,3}^{123},U_{1,2}^{12}),12$ Part 1 Thm.LABEL:part1-rk3
Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk1 $(U_{1,4},U^{12}_{1,2}),6$ Part 1 Thm.LABEL:part1-rk2 $(\mathcal{W}_{2}^{24},U_{1,2}^{14}),24$ $(U_{2,4},U^{12}_{1,2}),6$ $(U_{3,4},U^{12}_{1,2}),6$ $(\hat{U}^{1}_{2,5},U_{1,2}^{12}),12$ $(\hat{U}^{1}_{3,5},U^{12}_{1,2}),12$ $(V^{12}_{8},U^{13}_{1,2}),24$
$(U^{12}_{1,2},U^{34}_{1,2}),3$ $(U^{123}_{1,3},U^{14}_{1,2}),12$ Part 1 Thm.LABEL:part1-rk1 $(U_{2,3}^{123},U_{1,2}^{14}),12$ Part 1 Thm.LABEL:part1-rk3 Part 2 Thm.III-A.2 Part 1 Thm.LABEL:part1-rk9 Part 1 Thm.LABEL:part1-rk18 Part 1 Thm.LABEL:part1-rk21 Part 1 Thm.LABEL:part1-rk31
Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk3 $(\mathcal{W}_{2}^{34},U_{1,2}^{12}),6$
Part 1 Thm.LABEL:part1-rk2
[…]
Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk1 Part 1 Thm.LABEL:part1-rk4 Part 1 Thm.LABEL:part1-rk2 Part 2 Thm.III-A.1 Part 1 Thm.LABEL:part1-rk11 Part 1 Thm.LABEL:part1-rk19 Part 1 Thm.LABEL:part1-rk22 Part 1 Thm.LABEL:part1-rk31
$U_{1,4}$ $\backslash \\backslash \\$backslash $(U_{2,3}^{123},U_{1,4}),4$ 0 0 $(U_{3,4},U_{1,4}),1$ 0 0 $(V^{12}_{8},U_{1,4}),6$
Part 1 Thm.LABEL:part1-rk7 Part 1 Thm.LABEL:part1-rk12 Part 1 Thm.LABEL:part1-rk31
$U^{\beta}_{2,3}$ $\backslash \\backslash \\$backslash $(U_{2,3}^{123},U_{2,3}^{124}),6$ $(\mathcal{W}_{2}^{12},U_{2,3}^{134}),12$ $(U_{2,4},U^{123}_{2,3}),4$ $(U_{3,4},U^{123}_{2,3}),4$ $(\hat{U}^{1}_{2,5},U_{2,3}^{234}),4$ $(\hat{U}^{1}_{3,5},U^{123}_{2,3}),12$ $(V^{12}_{8},U^{123}_{2,3}),12$
Part 1 Thm.LABEL:part1-rk5 Part 1 Thm.LABEL:part1-rk6 Part 2 Thm.III-B.2 Part 1 Thm.LABEL:part1-rk10 Part 2 Thm.III-C.2 Part 1 Thm.LABEL:part1-rk23 Part 1 Thm.LABEL:part1-rk31
$\mathcal{W}^{\beta}_{2}$ $\backslash \\backslash \\$backslash $\backslash$ $(\mathcal{W}_{2}^{12},\mathcal{W}_{2}^{13}),12$ $(U_{2,4},\mathcal{W}^{12}_{2}),6$ 0 $(\hat{U}^{1}_{2,5},\mathcal{W}_{2}^{12}),12$ $(\hat{U}^{1}_{3,5},\mathcal{W}^{23}_{2}),12$ 0
Part 2 Thm.III-C.1 Part 2 Thm.III-B.1 Part 2 Thm.III-B.3 Part 1 Thm.LABEL:part1-rk24
$U_{2,4}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash 0 $(\hat{U}^{1}_{2,5},U_{2,4}),4$ $(\hat{U}^{1}_{3,5},U_{2,4}),4$ 0
Part 2 Thm.III-C.3 Part 2 Thm.III-D.1
$U_{3,4}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash $\backslash$ 0 0 $(V^{12}_{8},U_{3,4}),6$
Part 1 Thm.LABEL:part1-rk31
$\hat{U}^{j}_{2,5}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash $\backslash$ 0 0 0
$\hat{U}^{j}_{3,5}$ $\backslash \\backslash \\$$\backslash \\backslash \\$$\backslash \\backslash$ 0 0
$V^{\beta}_{8}$ $\backslash \\backslash \\$$\backslash \\backslash \\$$\backslash \\backslash \\$backslash 0
We label the rows and columns with the $11$ types of extreme rays of $\Gamma_{4}$. For simplicity, we denote the face $F=\mathrm{cone}(E_{1},E_{2})$ by $(E_{1},E_{2})$. In each cell with $(E_{1},E_{2})\ n$, Part $k$ Thm. mm”, $k=1$ or $2$, $(E_{1},E_{2})$” denote a representative of the type of $2$-dimensional faces, where $E_{1}$ $(E_{2})$” is a representative of the type the extrem rays in the column (row), “nn” is the number of the faces in this type and this face is characterized in “Thm. mm” of “Part kk” of this series. For the cell with “0”, the convex hull of the two extreme rays in each type forms no $2$-dimensional faces of $\Gamma_{4}$.
III Characterization of entropy functions on Two-dimensional faces of $\Gamma_{4}$
In this section, we characterize the entropy functions on $2$-dimensional faces of $\Gamma_{4}$. We embed each face $F=(E_{1},E_{2})$ in the first octant of a $2$-dimensional Cartesian coordinate system whose axes are labeled by $a$ and $b$. Thus, for each $(a,b)$, $a,b\geq 0$, it represents the polymatroid $a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{i},i=1,2$, is the rank function of the minimal integer polymatroid in $E_{i}$, respectively. Throughout this paper, for a random vector $(X_{i},i\in N_{4})$ or its subvectors, we assume each $X_{i}$ is distributed on a finite set $\mathcal{X}_{i}$, and for each $x_{i}\in X_{i}$, $p(x_{i})>0$.
Lemma 1.
([23, Lemma 1]) If $X_{1}$ and $X_{2}$ are independent and for any $p(x_{1},x_{2},x_{3})>0$, $p(x_{1})=p(x_{2})$, then $X_{1}$ and $X_{2}$ are uniformly distributed on $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, respectively, $|\mathcal{X}_{1}|=|\mathcal{X}_{2}|$ and $H(X_{1})=H(X_{2})$.
Lemma 2.
([1, Lemma 15.3]) For any $\mathbf{h}_{1},\mathbf{h}_{2}\in\Gamma^{*}_{n}$, $\mathbf{h}_{1}+\mathbf{h}_{2}\in\Gamma^{*}_{n}$.
III-A Entropy functions on faces with extreme rays containing $U_{2,4}$ and one rank-1 matroid
In this subsection, we characterize entropy functions on the $2$-dim faces of $\Gamma_{4}$ with extreme rays containing $U_{2,4}$ and one rank-1 matroid.
Theorem III-A.1.
For $F=(U_{2,4},U^{123}_{1,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if
  • $a+b\geq\log v$ and $\log(v-1)<a\leq\log v$ for positive integer $v\neq 2,6$; or
  • $a+b\geq\log{(v+1)}$ and $\log(v-1)<a\leq\log v$ for $v=2,6.$
[…]
Figure 1: The face $(U_{2,4},U_{1,3}^{123})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{i4})$ $=H(X_{i})+H(X_{4}),i\in\{1,2,3\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),|K|=2.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.1.1)
$=p(x_{1},x_{2},x_{4})$ (III.1.2)
$=p(x_{1},x_{3},x_{4})$ (III.1.3)
$=p(x_{2},x_{3},x_{4}),$ (III.1.4)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.1.5)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.1.6)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.1.7)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.1.8)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.1.9)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.1.10)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.1.11)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.1.12)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.1.13)
By (III.1.1)-(III.1.4) and (III.1.8)-(III.1.13), we have
$p(x_{1234})=p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})$ (III.1.14)
$=p(x_{1},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$ (III.1.15)
$=p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4})$ (III.1.16)
$=p(x_{2},x_{3})=p(x_{2},x_{4})=p(x_{3},x_{4}).$ (III.1.17)
Then by (III.1.5)-(III.1.7) and (III.1.16)-(III.1.17), we obtain
$p(x_{1})p(x_{4})=p(x_{2})p(x_{4})=p(x_{3})p(x_{4}).$ (III.1.18)
Canceling $p(x_{4})$ in the above equation,
$p(x_{1})=p(x_{2})=p(x_{3}).$ (III.1.19)
Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $\{1,2,4\}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $\{1,2,4\}$. It can be checked that they are the rank functions of $U_{2,3}$ and $U^{12,3}_{1,2}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12,3}_{1,2})$ and
$a+b\geq\log\lceil 2^{a}\rceil$ (III.1.20)
by [10, Theorem 1], which is an outer bound on the entropy region of $F$.
Now we show that
  • $a+b\geq\log v$ and $\log(v-1)<a\leq\log v$ for positive integer $v\neq 2,6$; or
  • $a+b\geq\log{(v+1)}$ and $\log(v-1)<a\leq\log v$ for $v=2,6$
form an inner bound on the entropy region on $F$. Let $\mathbf{T}$ be a $\mathrm{VOA}(U_{2,4},v)$. Let $X_{4}$ be distributed on $\mathbb{I}_{v}$ such that $H(X_{4})=a$. Let $(X_{i},i\in N_{4})$ be distributed on the rows of $\mathbf{T}$ with $p(x_{1},x_{2},x_{3},x_{4})=\frac{p(x_{4})}{v}$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,\log v-a)$. Then by Lemma 2, all $\mathbf{h}$ in the inner bound are entropic.
[…]
Figure 2: Inner and outer bounds on $F^{*}$
It can be seen that there exists a gap between the inner and outer bounds,
$\log{v}\leq a+b<\log{(v+1)}$ (III.1.21)
and
$\log{(v-1)}<a\leq\log{v}\text{ for }v=2,6,$ (III.1.22)
which is the slash region in Fig. 2. In the following we prove that polymatroids in this gap are all non-entropic.
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. By $p(x_{1},x_{2},x_{4})=p(x_{1},x_{2})$, each edge can be colored by a unique $x_{4}$. By $p(x_{1},x_{2},x_{4})=p(x_{1},x_{4})$, any two edges incident to $x_{1}$ are colored differently. Due to (III.1.5), $X_{1}$ and $X_{4}$ are independent, thus all colors will occur at least once on the edges incident to $x_{1}$. Hence, for each vertex $x_{1}$, it is incident to $k$ edges where $v=|\mathcal{X}_{4}|$. It holds for each $x_{2}$ as well by symmetry. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$, that is, the probability of the event that the random vector takes a tuple in $C_{j}$. In a connected component $C_{j}$, the number of edges is $n_{1}^{(j)}v=n_{2}^{(j)}v$, which implies that $n_{1}^{(j)}=n_{2}^{(j)}$, and so it can be simplified to $n^{(j)}$. Since each vertex is incident to $v$ edges, we have $n^{(j)}\geq v$. As $p(x_{1},x_{2})$ $\sum_{x_{3},x_{4}}p(x_{1},x_{2},x_{3},x_{4})$, there exist $x_{3},x_{4}$ such that $p(x_{1},x_{2},x_{3},x_{4})>0$. By (III.1.19), $p(x_{1})=p(x_{2})$, which implies that the probability mass of the two adjacent vertices are the same, and so are the vertices in a connected compoent as well. Since $p(x_{1},x_{2},x_{3})=p(x_{1},x_{2})$, we color each $(x_{1},x_{2})$ by $x_{3}\in\mathcal{X}_{3}$. As $p(x_{1})=p(x_{2})=p(x_{3})$ by (III.1.19), the probability mass of the color $x_{3}$ in a connected component in $G$ are the same. We classify the connected components of $G$ into $t_{1}$ equivalence class such that for any two components in an equivalence class, they share a common color in $\mathcal{X}_{3}$. Let $A_{i},i=1,2,\dots,t_{1}$ be index set of the components in each equivalence class $i$. Thus, the probability mass of vertices in the same equivalence class in $G$ are equal. Then
$H(X_{1})$ $=-\sum_{j=1}^{t_{1}}n^{(j)^{\prime}}\dfrac{p^{\prime}_{j}}{n^{(j)^{\prime}}}\log\dfrac{p^{\prime}_{j}}{n^{(j)^{\prime}}}$ (III.1.23)
$=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}},$ (III.1.24)
where $p^{\prime}_{j}\triangleq\sum_{i\in A_{j}}p_{i}$, $n^{(j)^{\prime}}\triangleq\sum_{i\in A_{j}}n^{(i)}$.
As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=H(X_{2})=H(X_{3})$ (III.1.25)
$=a+b,$ (III.1.26)
$H(X_{4})$ $=a.$ (III.1.27)
By (III.1.24) and $n^{(j)}\geq v$,
$a+b$ $=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}}$ (III.1.28)
$\geq H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log v$ (III.1.29)
$\geq\log v.$ (III.1.30)
Note that $|\mathcal{X}_{4}|=v$, we have $a\leq\log v$. Then
$a+b\geq\log v\geq a.$ (III.1.31)
Assume there exists one equivalence class that contains only one connected component $C_{j}$ and satisfies $n^{(j)}=v$ for $v=2,6$. Then for the connected component $C_{j}$, as $p(x_{1})=p(x_{2})=p(x_{3})$,
$n_{3}^{(j)}=\dfrac{p^{\prime}_{j}}{p(x_{3})}=\dfrac{p_{j}}{p(x_{1})}=n^{(j)},$ (III.1.32)
where $n_{3}^{(j)}$ denotes the number of the colors $x_{3}$ in $\mathcal{X}_{3}$ in $C_{j}$. Let $\mathbf{T}$ be a $v^{2}\times 4$ array, and for each row of $\mathbf{T}$, the four entries correspond to the two ends of an edge $(x_{1},x_{2})$ in $C_{j}$, and the color in $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$ of the edge, respectively. It is easy to check that both $\mathbf{T}(1,2,3)$ and $\mathbf{T}(1,2,4)$ are $\mathrm{VOA}(U_{2,3},v)$s. Since $X_{3}$ and $X_{4}$ are independent, each pair $(x_{3},x_{4})\in\mathbb{I}^{2}_{v}$ appears in $\mathbf{T}(3,4)$ as a row exactly once. Therefore, $\mathbf{T}$ is a $\mathrm{VOA}(U_{2,4},v)$, which contradicts the non-existence of $\mathrm{VOA}(U_{2,4},2)$ or $\mathrm{VOA}(U_{2,3},6)$. Hence, each equivalence class either contains multiple connected components or contains only one connected component satisfying $n^{(j)}>v$. Thus,
$n^{(j)^{\prime}}=\sum_{i\in A_{j}}n^{(i)}\geq v+1.$ (III.1.33)
By (III.1.24) and (III.1.26),
$a+b$ $=H(X_{1})=H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{n^{(j)^{\prime}}}$ (III.1.34)
$\geq H(p^{\prime}_{1},\dots,p^{\prime}_{t_{1}})+\sum_{j=1}^{t_{1}}p^{\prime}_{j}\log{(v+1)}$ (III.1.35)
$\geq\log{(v+1)},$ (III.1.36)
which implies the gap is non-entropic. ∎
Theorem III-A.2.
For $F=(U_{2,4},U^{12}_{1,2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if
  • $a=\log v$ for positive integer $v\neq 2,6$;
  • $a=\log 2$, $b\geq\log 2$; or
  • $a=\log 6$, $b\geq\log 2$.
Figure 3: The face $(U_{2,4},U^{12}_{1,2})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),\{i,j\}\neq\{1,2\}$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad|K|=2,K\neq\{3,4\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.1.37)
$=p(x_{1},x_{2},x_{4})$ (III.1.38)
$=p(x_{1},x_{3},x_{4})$ (III.1.39)
$=p(x_{2},x_{3},x_{4}),$ (III.1.40)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (III.1.41)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.1.42)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.1.43)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.1.44)
$p(x_{3},x_{4})$ $=p(x_{4})p(x_{4}),$ (III.1.45)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.1.46)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.1.47)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (III.1.48)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.1.49)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}).$ (III.1.50)
According to (III.1.37), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.1.46), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.1.51)
Together with (III.1.38), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (III.1.52)
By the same argument, we have
$p(x_{1234})$ $=p(x_{1},x_{3})=p(x_{1},x_{4})$ (III.1.53)
$=p(x_{2},x_{3})=p(x_{2},x_{4}).$ (III.1.54)
By (III.1.41) and (III.1.42), replacing $p(x_{1},x_{3})$ and $p(x_{1},x_{4})$ by $p(x_{1})p(x_{3})$ and $p(x_{1})p(x_{4})$ in (III.1.53), we obtain
$p(x_{3})=p(x_{4}).$ (III.1.55)
Since $X_{3}$ and $X_{4}$ are independent, by Lemma 1, $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and $H(X_{3})=H(X_{4})=\log v$ where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{3})$ $=H(X_{4})=a,$ (III.1.56)
which implies that $a$ can only take the value of $\log v$. By Lemma 2 and the fact that $a=\log v$ for $v\neq 2,6$ on the ray $U_{2,4}$ and the whole ray $U^{12}_{1,2}$ are entropic, all $\mathbf{h}=(a,b)\in F$ are entropic when $a=\log v$ for positive integer $v\neq 2,6$, and $b\geq 0$.
Now, we show that when $a=\log 2$ or $\log 6$,
$b\geq\log 2.$ (III.1.57)
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. Assume $G$ has $t$ connected components and $|\mathcal{X}_{3}|=|\mathcal{X}_{4}|=v$ for $v=2,6$. By (III.1.37), (III.1.38) and (III.1.51), we have $p(x_{1},x_{2})=p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})$, which implies that $X_{i}$ is a function of $X_{1}$ and $X_{2}$, $i=3,4$. Then each edge $(x_{1},x_{2})$ can be colored by a unique $x_{3}\in\mathcal{X}_{3}$ and a unique $x_{4}\in\mathcal{X}_{4}$. By (III.1.37) and (III.1.53), we have $p(x_{1},x_{3})=p(x_{1},x_{2},x_{3})$. Thus, any two edges incident to $x_{1}$ are colored by different $x_{3}\in\mathcal{X}_{3}$. Since $X_{1}$ and $X_{3}$ are independent, all colors $x_{3}\in\mathcal{X}_{3}$ will occur at least once on the edges adjacent to $x_{1}$. Hence, each $x_{1}$ is incident to $k$ edges. It holds for each $x_{2}$ as well by symmetry. We denote the number of the vertices in $\mathcal{X}_{i}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ in the connected component $C_{j}$ and the probability mass of $C_{j}$ by $p_{j}$. In each connected component $C_{j}$, the number of edges is $n_{1}^{(j)}k=n_{2}^{(j)}k$, which implies $n_{1}^{(j)}=n_{2}^{(j)}$ and so it can be simplified to $n^{(j)}$. As $p(x_{1},x_{2})=\sum_{x_{3},x_{4}}p(x_{1},x_{2},x_{3},x_{4})$, there exist $x_{3},x_{4}$ such that $p(x_{1},x_{2},x_{3},x_{4})>0$. Due to (III.1.53) and (III.1.54), we have $p(x_{1},x_{3})=p(x_{2},x_{3})$. By (III.1.41) and (III.1.43), replacing $p(x_{1},x_{3})$ and $p(x_{2},x_{3})$ by $p(x_{1})p(x_{3})$ and $p(x_{2})p(x_{3})$, we obtain
$p(x_{1})=p(x_{2}),$ (III.1.58)
which implies the probability mass of two adjacent vertices are the same, and so are the vertices in a connected compoent as well. In each connected component $C_{j}$, the probability mass of the vertices are equal to
$p(x_{1})=p(x_{2})=\dfrac{p_{j}}{n^{(j)}}.$ (III.1.59)
By (III.1.41), (III.1.52) and (III.1.53), the probability mass of the edges are equal to
$p(x_{12})$ $=p(x_{1234})=p(x_{1},x_{3})=p(x_{1})p(x_{3}).$ (III.1.60)
Note that $X_{3}$ is uniformly distributed on $\mathcal{X}_{3}$ and $|\mathcal{X}_{3}|=v$. Replacing $p(x_{3})$ by $\frac{1}{v}$ in (III.1.60), we have
$p(x_{1},x_{2})=p(x_{1})p(x_{3})=\dfrac{p(x_{1})}{v}=\dfrac{p_{j}}{n^{(j)}v}.$ (III.1.61)
Hence,
$H(X_{1})$ $=-\sum_{j=1}^{t}n^{(j)}\dfrac{p_{j}}{n^{(j)}}\log\dfrac{p_{j}}{n^{(j)}}$ (III.1.62)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}.$ (III.1.63)
$H(X_{1},X_{2})=-\sum_{j=1}^{t}n^{(j)}v\dfrac{p_{j}}{n^{(j)}v}\log\dfrac{p_{j}}{n^{(j)}v}$ (III.1.64)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}v$ (III.1.65)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}+\sum_{j=1}^{t}p_{j}\log v$ (III.1.66)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}+\log v.$ (III.1.67)
As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=a+b,$ (III.1.68)
$H(X_{1},X_{2})$ $=2a+b.$ (III.1.69)
Equating the above equations with (III.1.63) and (III.1.67), we obtain
$a$ $=\log v,$ (III.1.70)
$b$ $=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}-\log v.$ (III.1.71)
Now we show that for each connected component, there exists an $(x_{3},x_{4})$ that occurs at least twice. By (III.1.45), $X_{3}$ and $X_{4}$ are independent. Since $|\mathcal{X}_{3}|=|\mathcal{X}_{4}|=v$, the number of the pairs $(x_{3},x_{4})$ is $v^{2}$. The probability mass of each $(x_{3},x_{4})$ is
$p(x_{3},x_{4})=p(x_{3})p(x_{4})=\dfrac{1}{v^{2}}.$ (III.1.72)
In each connected component $C_{j}$, when $n^{(j)}>v$, there are $n^{(j)}v$ edges. Due to the pigeonhole principle, there exists an $(x_{3},x_{4})$ that occurs at least twice. When $n^{(j)}=v$, let $\mathbf{T}$ be a $v^{2}\times 4$ array, and for each row of $\mathbf{T}$, the four entries correspond to the two ends of an edge $(x_{1},x_{2})$ in $C_{j}$, and the colors in $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$ of the edge, respectively. It is easy to check that both $\mathbf{T}(1,2,3)$ and $\mathbf{T}(1,2,4)$ are $\mathrm{VOA}(U_{2,3},v)$s. The non-existence of $\mathrm{VOA}(U_{2,4},2)$ or $\mathrm{VOA}(U_{2,4},6)$ implies that the number of different entries $(x_{3},x_{4})$ on the rows of $\mathbf{T}(3,4)$ is less than $v^{2}$, so there exists an $(x_{3},x_{4})$ that occurs at least twice.
The probability mass of the pair $(x_{3},x_{4})$ that occurs at least twice in $C_{j}$ is
$p(x_{3},x_{4})=\sum_{x_{1},x_{2}}p(x_{1},x_{2},x_{3},x_{4}).$ (III.1.73)
By (III.1.60) and (III.1.61),
$p(x_{1},x_{2},x_{3},x_{4})=\dfrac{p_{j}}{n^{(j)}v},$ (III.1.74)
which is independent of $(x_{1},x_{2})$ and so
$p(x_{3},x_{4})$ $=\sum_{x_{1},x_{2}}p(x_{1},x_{2},x_{3},x_{4})\geq\dfrac{2p_{j}}{n^{(j)}v}.$ (III.1.75)
Together with (III.1.72), we obtain
$\dfrac{1}{v^{2}}\geq\dfrac{2p_{j}}{n^{(j)}v},$ (III.1.76)
which implies
$n^{(j)}\geq 2vp_{j}.$ (III.1.77)
Substituting (III.1.77) into (III.1.71),
$b$ $=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)}-\log v$ (III.1.78)
$\geq H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log 2vp_{j}-\log v$ (III.1.79)
$=H(p_{1},...,p_{t})-H(p_{1},...,p_{t})+\log 2v-\log v$ (III.1.80)
$=\log 2.$ (III.1.81)
To prove the theorem, it remains to show that all $(a,b)\in F$ are entropic if $a=\log 2$ or $\log 6$, $b\geq\log 2$, which can be implied by the fact that polymatroids $\mathbf{h}=(\log 2,\log 2)$ and $(\log 6,\log 2)$ are entropic. For $v=2,6$, let $\mathbf{T}_{v}$ be a $2v^{2}\times 4$ array with entries in $\mathbb{I}_{2v}$ such that
  • each pair in $\mathbb{I}^{2}_{v}\cup\mathbb{I}^{\prime 2}_{v}$ occurs exactly once in $\mathbf{T}_{v}(1,2)$, where $\mathbb{I}^{\prime}_{v}=\mathbb{I}_{2v}\setminus\mathbb{I}_{v}$;
  • each pair in $\mathbb{I}^{2}_{v}$ occurs exactly twice in $\mathbf{T}_{v}(3,4)$; and
  • $\mathbf{T}^{\prime}_{v}(1,2,3)$, $\mathbf{T}^{\prime}_{v}(1,2,4)$, $\mathbf{T}^{\prime\prime}_{v}(1,2,3)$ and $\mathbf{T}^{\prime\prime}_{v}(1,2,4)$ are all $\mathrm{VOA}(U_{2,3},v)$s, where $\mathbf{T}^{\prime}_{v}$ is a $v^{2}\times 4$ subarray of $\mathbf{T}_{v}$ formed by the rows with first two entries in $\mathbb{I}_{v}$, and $\mathbf{T}^{\prime\prime}_{v}$ is a $v^{2}\times 4$ subarray of $\mathbf{T}_{v}$ formed by the rows with first two entries in $\mathbb{I}^{\prime}_{v}$.
It can be seen that the following $\mathbf{T}_{2}$ and $\mathbf{T}_{6}$ are such constructed.
$\mathbf{T}_{2}\ =\ \begin{matrix}0&0&0&0\\ 0&1&1&1\\ 1&0&1&1\\ 1&1&0&0\\ 2&2&0&1\\ 2&3&1&0\\ 3&2&1&0\\ 3&3&0&1\end{matrix}$
[…]
Due to page limitation, disjoint subarrays $\mathbf{T}^{\prime}_{6}$ and $\mathbf{T}^{\prime\prime}_{6}$ of $\mathbf{T}_{6}$ are juxtaposed. Let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}_{v}$, $v=2,6$. It can be checked that for any nonempty $A\subseteq N_{4}$,
[…]
It can be checked that the entropy function is in $F$. Then by (III.1.68), (III.1.69) and (III.1.82), we have $a=\log v$ and $b=\log 2$. The proof is accomplished. ∎
Theorem III-A.3.
For $F=(U_{2,4},U^{4}_{1,1})$, $\mathbf{h}=(a,b)\in F$ is
  • entropic if $a=\log v$ for integer $v\neq 2,6$ or $a=\log 6,b\geq\log 2$; and
  • non-entropic if $a\neq\log v$ for some integer $v>0$ or $a=\log 2$.
Figure 4: The face $(U_{2,4},U^{4}_{1,1})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in\{1,2,3\},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in N_{4},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\quad|K|=2,K\subseteq N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{4})$ (III.1.83)
$=p(x_{1},x_{3},x_{4})$ (III.1.84)
$=p(x_{2},x_{3},x_{4}),$ (III.1.85)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (III.1.86)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (III.1.87)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.1.88)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.1.89)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.1.90)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.1.91)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.1.92)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.1.93)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234})$ (III.1.94)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.1.95)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.1.96)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.1.97)
By (III.1.83), canceling $p(x_{1},x_{2},x_{4})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.1.94), we obtain
$p(x_{1},x_{3},x_{4})=p(x_{1},x_{4}).$ (III.1.98)
Together with (III.1.84), we have
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{4})$ (III.1.99)
By the same argument, we obtain
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4}),$ (III.1.100)
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4}).$ (III.1.101)
Restricting $\mathbf{h}$ on $M=\{1,2,3\}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}$, where $\mathbf{r}^{\prime}$ is the rank function of $U_{2,3}$ on $M$. Thus the characterizing random vector $(X_{1},X_{2},X_{3})$ of $\mathbf{h}^{\prime}$ is uniformly distributed on the rows of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$ for a positive integer $v$, and so $a$ can only take the value of $\log v$. Note that $a=\log v$, $v\neq 2,6$, on the ray $U_{2,4}$, and the whole ray $U_{1,1}^{4}$ are entropic. By Lemma 2, $\mathbf{h}=(\log v,b)$ is entropic for positive integers $v$ with $v\neq 2,6$ and $b\geq 0$.
Now we only need to consider $\mathbf{h}=(\log v,b)$ for $v=2,6$, $b\geq 0$. Assume $\mathbf{h}=(\log 2,b)$ is entropic. Note that up to isomorphism, there exists only one $\mathrm{VOA}(U_{2,3},2)$. Without loss of generality, let the characterizing random vector $(X_{1},X_{2},X_{3})$ of $\mathbf{h}^{\prime}$ be uniformly distributed on the rows of $\mathbf{T}$ as follows:
$\mathbf{T}\ =\ \begin{matrix}0&0&0\\ 0&1&1\\ 1&0&1\\ 1&1&0\end{matrix}$
As $p_{X_{1}X_{2}X_{3}}(0,0,0)>0$, there exists $x_{4}\in\mathcal{X}_{4}$ such that $p_{X_{1}X_{2}X_{3}X_{4}}(0,0,0,x_{4})>0$. Note that $X_{1}$ and $X_{4}$ are independent by (III.1.88), we obtain
$p_{X_{1}X_{4}}(1,x_{4})=p_{X_{1}}(1)p_{X_{4}}(x_{4})>0,$ (III.1.102)
which implies that either $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})>0$ or $p_{X_{1}X_{2}X_{3}X_{4}}(1,1,0,x_{4})>0$. Since $p_{X_{1}X_{2}X_{3}X_{4}}(x_{1},x_{2},x_{3},x_{4})=p_{X_{2}X_{4}}(x_{2},x_{4})$ by (III.1.100),
$p_{X_{1}X_{2}X_{3}X_{4}}(0,0,0,x_{4})=p_{X_{2}X_{4}}(0,x_{4})$ (III.1.103)
However,
$p_{X_{2}X_{4}}(0,x_{4})=\sum_{x_{1},x_{3}}p_{X_{1}X_{2}X_{3}X_{4}}(x_{1},0,x_{3},x_{4})$ (III.1.104)
$=p_{X_{1}X_{2}X_{3}X_{4}}(0,0,0,x_{4})+p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4}),$ (III.1.105)
which implies $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})=0$ contradicting $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})>0$. Similarly, we can show that by (III.1.101), $p_{X_{1}X_{2}X_{3}X_{4}}(1,1,0,x_{4})>0$ will also lead a contradiction.
As for $\mathbf{h}=(\log 6,b)$, we will show an inner bound on the entropy region within these polymatroids, i.e, those with $b\geq 0$. Let $\mathbf{T}^{\prime}$ be the array as follows.
Written down within a page, the first to $36$th and $37$th to $72$th rows of $\mathbf{T^{\prime}}$ are juxtaposed. Let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}^{\prime}$. Then it can be checked that such construction is $(\log 6,\log 2)$. Then by the fact that the whole ray $U^{4}_{1,1}$ are entropic and Lemma 2, all $\mathbf{h}=(\log 6,b)$ with $b\geq\log 2$ are entropic. ∎
Remark In this theorem, we give an inner bound on the face $(U_{2,4},U^{4}_{1,1})$. Entropy functions on this face corresponds to a pair orthogonal squares, one is Latin and the other is multi-symbol Latin. A square is called a multi-symbol Lain square of order $v$ with symbol set size $v^{\prime}\geq v$ if it is a $v\times v$ square with a set of symbols with size $v^{\prime}$ and each cell contains one or more symbols, and each symbol appears in each row and each column exactly once. Such kind of pair of squares can be obtained by splitting the symbols of one square of a pair of orthogonal Latin squares for $v\neq 2,6$. We proved that such a pair does not exist for $v=2$. For $v=6$, we gave a pair with $v^{\prime}=12$. We conjecture that this inner bound is tight.
III-B Entropy functions on faces with both extreme rays containing rank $2$ integer polymatroids
In this subsection, we characterize entropy functions on three $2$-dimensional faces $(U_{2,4},\mathcal{W}^{12}_{2})$, $(U_{2,4},U^{123}_{2,3})$, and $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$ of $\Gamma_{4}$ with extreme rays both containing rank $2$ integer polymatroids. Some other faces in this family $(U_{2,3}^{123},U_{2,3}^{124})$ and $(\mathcal{W}_{2}^{12},U_{2,3}^{134})$ have already been characterized in Part I of this series of two papers, while $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$ will be characterized in Subsection III-C as Latin square decomposition will be used.
$\mathbf{T}^{\prime}\ =\ \begin{matrix}0&0&0&0\\ 0&1&5&7\\ 0&2&3&2\\ 0&3&4&1\\ 0&4&2&4\\ 0&5&1&5\\ 1&0&1&8\\ 1&1&0&2\\ 1&2&4&5\\ 1&3&3&4\\ 1&4&5&0\\ 1&5&2&7\\ 2&0&2&2\\ 2&1&4&6\\ 2&2&1&0\\ 2&3&5&9\\ 2&4&3&5\\ 2&5&0&4\\ 3&0&3&1\\ 3&1&2&5\\ 3&2&5&4\\ 3&3&1&2\\ 3&4&0&6\\ 3&5&4&0\\ 4&0&4&4\\ 4&1&3&0\\ 4&2&2&1\\ 4&3&0&5\\ 4&4&1&3\\ 4&5&5&2\\ 5&0&5&5\\ 5&1&1&4\\ 5&2&0&7\\ 5&3&2&0\\ 5&4&4&2\\ 5&5&3&3\end{matrix}\qquad\qquad\begin{matrix}0&0&0&9\\ 0&1&5&11\\ 0&2&3&8\\ 0&3&4&3\\ 0&4&2&10\\ 0&5&1&6\\ 1&0&1&10\\ 1&1&0&3\\ 1&2&4&9\\ 1&3&3&6\\ 1&4&5&1\\ 1&5&2&11\\ 2&0&2&3\\ 2&1&4&8\\ 2&2&1&11\\ 2&3&5&10\\ 2&4&3&7\\ 2&5&0&1\\ 3&0&3&11\\ 3&1&2&9\\ 3&2&5&3\\ 3&3&1&7\\ 3&4&0&8\\ 3&5&4&10\\ 4&0&4&7\\ 4&1&3&10\\ 4&2&2&6\\ 4&3&0&11\\ 4&4&1&9\\ 4&5&5&8\\ 5&0&5&6\\ 5&1&1&1\\ 5&2&0&10\\ 5&3&2&8\\ 5&4&4&11\\ 5&5&3&9\end{matrix}$
Theorem III-B.1.
For $F=(U_{2,4},\mathcal{W}^{12}_{2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log v$ for integer $v>0$, and there exists a $v^{2}\times 4$ array $\mathbf{T}$ such that $\mathbf{T}(1,3,4)$ and $\mathbf{T}(2,3,4)$ are $\mathrm{VOA}(U_{2,3},v)$, and
$a=H(\bm{\alpha})-\log v,$
where $\bm{\alpha}=(\alpha_{x_{1},x_{2}}>0:x_{1},x_{2}\in\mathbb{I}_{v})$ and $\alpha_{x_{1},x_{2}}$ denotes the times of the row $(x_{1},x_{2})$ that occurs in $\mathbf{T}(1,2)$.
[…]
Figure 5: The face $(U_{2,4},W^{12}_{2})$
Proof.
Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $M=\{1,3,4\}$ or $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=(a+b)\mathbf{r}^{\prime}$, where $\mathbf{r}^{\prime}$ is the rank function of $U_{2,3}$ on $M$. As $\mathbf{X}=(X_{i},i\in N_{4})$ is the characterizing random vector of $\mathbf{h}$, $(X_{1},X_{3},X_{4})$ and $(X_{2},X_{3},X_{4})$ are uniformly distributed on the rows of some $\mathrm{VOA}(U_{2,3},v)s$, which implies that $a+b=\log v$ where $v=|\mathcal{X}_{1}|=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. Let $\mathbf{T}$ be a $v^{2}\times 4$ array such that both $\mathbf{T}(1,3,4)$ and $\mathbf{T}(2,3,4)$ are $\mathrm{VOA}(U_{2,3},v)s$. If $\mathbf{h}\in F$ is entropic, $\mathbf{X}$ satisfies the following information equalities
$H(X_{1}|X_{2},X_{3},X_{4})=H(X_{2}|X_{1},X_{3},X_{4})=0,$ (III.2.1)
which implies that $\mathbf{X}$ must be uniformly distributed on such a constructed $\mathbf{T}$. Thus the probability mass of each pair $(x_{1},x_{2})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$ is
$p(x_{1},x_{2})=\sum_{x_{3},x_{4}:\atop p(x_{1},x_{2},x_{3},x_{4})>0}p(x_{1},x_{2},x_{3},x_{4})=\frac{\alpha_{x_{1},x_{2}}}{v^{2}},$ (III.2.2)
where $\alpha_{x_{1},x_{2}}$ denotes the times of the row $(x_{1},x_{2})$ that occurs in $\mathbf{T}(1,2)$. Hence, $\bm{\alpha}=(\alpha_{x_{1},x_{2}}>0:x_{1},x_{2}\in\mathbb{I}_{v})$ forms a partition of $v^{2}$. Then
$H(X_{1},X_{2})=H(\bm{\alpha}).$ (III.2.3)
As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=a+b=\log v,$ (III.2.4)
$H(X_{1},X_{2})$ $=2a+b.$ (III.2.5)
By (III.2.3)-(III.2.5), we conclude that
$a=H(\bm{\alpha})-\log v.$ (III.2.6)
Now we prove the “if ” part. Since there exists a $v^{2}\times 4$ array $\mathbf{T}$ satisfying the sufficient condition, let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}$. It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$. ∎
Remark This theorem studies the relations of two Latin squares (of first order) of the same size from the perspective of entropy functions, which generalize both orthogonal Latin squares and two identical Latin squares. For any pair of two $v\times v$ Latin squares $S_{1}$ and $S_{2}$, it correspond to a $v^{2}\times 4$ $\mathbf{T}$ array with each row $(i,j,s_{1},s_{2})$, where $i,j\in\mathbb{I}_{v}$ and $s_{1}$ and $s_{2}$ are the symbols in $S_{1}(i,j)$ and $S_{2}(i,j)$, respectively. On one hand, when $S_{1}$ and $S_{2}$ are orthogonal, entropy function $\mathbf{h}$ of a random vector distributed on the rows of $\mathbf{T}$ is in the extreme ray containing $U_{2,4}$; on the other hand, when they are identical, $\mathbf{h}$ is in the extreme ray containing $\mathcal{W}^{12}_{2}$. When they are neither orthogonal nor identical, $\mathbf{h}$ locates in the face, but not in the extreme ray.
Theorem III-B.2.
For $F=(U_{2,4},U^{123}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log{v},a=H(\bm{\alpha})$ and $(a,b)\neq(\log 2,0),(\log 6,0)$, where integer $v>0$ and $\bm{\alpha}$ is a partition of $v$.
[…]
Figure 6: The face $(U_{2,4},U^{123}_{2,3})$
Proof.
For entropic $\mathbf{h}\in F$, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in N_{4},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\quad|K|=2,K\subseteq\{1,2,3\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.2.7)
$=p(x_{1},x_{2},x_{4})$ (III.2.8)
$=p(x_{1},x_{3},x_{4})$ (III.2.9)
$=p(x_{2},x_{3},x_{4}),$ (III.2.10)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (III.2.11)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (III.2.12)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.2.13)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.2.14)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.2.15)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.2.16)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.2.17)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.2.18)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}).$ (III.2.19)
By (III.2.7), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on both side of (III.2.17), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.2.20)
Equating (III.2.11) and (III.2.20) implies
$p(x_{1},x_{2},x_{4})=p(x_{1})p(x_{2}).$ (III.2.21)
By the same argument,
$p(x_{1},x_{3},x_{4})=p(x_{1})p(x_{3}),$ (III.2.22)
$p(x_{2},x_{3},x_{4})=p(x_{2})p(x_{3}).$ (III.2.23)
By (III.2.8)-(III.2.10), $p(x_{1},x_{2},x_{4})=p(x_{1},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$, together with (III.2.21)-(III.2.23), we have
$p(x_{1})=p(x_{2})=p(x_{3}).$ (III.2.24)
By (III.2.11), (III.2.12) and (III.2.14), $X_{1}$, $X_{2}$ and $X_{3}$ are pairwise independent. By Lemma 1, $X_{i}$ are uniformly distributed on $\mathcal{X}_{i}$ for $i=1,2,3$, and
$H(X_{1})=H(X_{2})=H(X_{3})=\log v,$ (III.2.25)
where $v=|\mathcal{X}_{1}|=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$. As $\mathbf{h}\in F$, $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=a+b,$ (III.2.26)
$H(X_{4})$ $=a,$ (III.2.27)
which implies that $a+b=\log v$. Note that $X_{1}$ and $X_{2}$ are uniformly distributed, then $p(x_{1})=p(x_{2})=\frac{1}{v}$. By (III.2.8) and (III.2.21), we have
$p(x_{1234})=p(x_{1},x_{2},x_{4})=p(x_{1})p(x_{2})=\frac{1}{v^{2}}.$ (III.2.28)
As
$p(x_{1},x_{4})=\sum_{x^{\prime}_{2},x^{\prime}_{3}:\atop p(x_{1},x^{\prime}_{2},x^{\prime}_{3},x_{4})>0}p(x_{1},x^{\prime}_{2},x^{\prime}_{3},x_{4}),$ (III.2.29)
and the choice of $(x_{2},x_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}$ can be arbitrary for a fixed $(x_{1},x_{4})\in\mathcal{X}_{1}\times\mathcal{X}_{4}$, replacing $p(x_{1},x^{\prime}_{2},x^{\prime}_{3},x_{4})$ by $\frac{1}{v^{2}}$ in (III.2.29), we obtain
$p(x_{1},x_{4})=\sum_{x^{\prime}_{2},x^{\prime}_{3}:\atop p(x_{1},x^{\prime}_{2},x^{\prime}_{3},x_{4})>0}\frac{1}{v^{2}}=\frac{\alpha(x_{1},x_{4})}{v^{2}},$ (III.2.30)
where $\alpha(x_{1},x_{4})\triangleq|\{(x^{\prime}_{2},x^{\prime}_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}:p(x_{1},x’_{2},x^{\prime}_{3},x_{4})>0\}|$. By (III.2.13), we have
$p(x_{1},x_{4})=p(x_{1})p(x_{4})=\frac{1}{v}p(x_{4}).$ (III.2.31)
In light of (III.2.30) and (III.2.31), we obtain
$p(x_{4})=\frac{\alpha(x_{4})}{v}$ (III.2.32)
where $\alpha(x_{4})=\alpha(x_{1},x_{4})$ for any $x_{1}\in\mathcal{X}_{1}$. Together with (III.2.27), it can be seen that
$a=H(X_{4})=H(\bm{\alpha}),$ (III.2.33)
where $\bm{\alpha}\triangleq(\alpha(x_{4}),x_{4}\in\mathcal{X}_{4})$ is a number partition of $v$.
So far, we have proved that
$\{\mathbf{h}=(a,b)\in F:\$$a+b=\log v,\quad v\in\mathbb{Z}^{+}$
$a=H(\bm{\alpha}),\quad\bm{\alpha}\text{ is a partition of }v\}$
forms an outer bound on the entropic region in $F=(U_{2,4},U^{123}_{2,3})$. To prove the theorem, we now only need to check whether this outer bound is tight. We will see in the following that all $\mathbf{h}$ in it are entropic except for $(\log 2,0)$ and $(\log 6,0)$.
  • For any positive integer $v\neq 2,6$, let $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{t})$ be a partition of $v$, and $\{A_{i},i=1,\dots,t\}$ be a partition of $\mathbb{I}_{v}$ with $|A(i)|=\alpha_{i}$. Let $\mathbf{T}$ be a $\mathrm{VOA}(U_{2,4},v)$, and $\mathbf{T^{\prime}}$ be a $v^{2}\times 4$ array such that $\mathbf{T}^{\prime}(i)=\mathbf{T}(i)$ for $i=1,2,3$ and each entry in $\mathbf{T}^{\prime}(4)$ be $j$ if the corresponding entry in $\mathbf{T}(4)$ is in $A_{j}$. Then let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T^{\prime}}.$ It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$.
    Example 3.
    Let $\mathbf{h}=(a,b)\in F$ with $a+b=\log 3$ and $a=H((1,2))$. Let
[…]
    Note that $\mathbf{T}$ is a $\mathrm{VOA}(U_{2,4},3)$ and $\bm{T}^{\prime}$ is constructed as above with $A_{1}=\{0\}$ and $A_{2}=\{1,2\}$. Let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T^{\prime}}.$ We can see that $H(X_{4})=H((1,2))=a$, and $\mathbf{h}$ is the entropy function of $(X_{i},i\in N_{4})$.
For $v=2$ and $6$, as there is no such $\mathrm{VOA}(U_{2,4},2)$ or $\mathrm{VOA}(U_{2,4},6)$ [7, Proposition 2], the above construction is invalid, and so we have to discuss them separately.
  • For $v=2$, there are only two partitions $(2)$ and $(1,1)$, which correspond to $(0,\log 2)$ and $(\log 2,0)$, respectively. For $(0,\log 2)$, it is entropic as $\mathrm{VOA}(U_{2,3},2)$ is constructible, while $(\log 2,0)$ is non-entropic as $\mathrm{VOA}(U_{2,4},2)$ is not constructible.
  • For $v=6$, let
[…]
    Note that both $\mathbf{T}_{\rm Eu}(\{1,2,3\})$ and $\mathbf{T}_{\rm Eu}(\{1,2,4\})$ are $\mathrm{VOA}(U_{2,3},6)$s. However, $\mathbf{T}_{\rm Eu}$ is not a $\mathrm{VOA}(U_{2,4},6)$, as only $34$ different pairs occur in $\mathbf{T}_{\rm Eu}(\{2,4\})$, while $(1,5)$ and $(3,4)$ each appear twice in $\mathbf{T}_{\rm Eu}(\{3,4\})$.  [^4] The array $\mathbf{T}_{\rm Eu}$ is constructed from the following two Latin squares
    0 1 2 3 4 5 1 2 5 4 0 3 2 3 0 1 5 4 3 5 4 2 1 0 4 0 1 5 3 2 5 4 3 0 2 1
    0 5 3 4 2 1 1 0 4 3 5 2 2 4 1 5 3 0 3 2 5 1 0 4 4 3 2 0 1 5 5 1 0 2 4 3
    discovered by Euler in [31] in 1782, in a manner how we construct $\mathrm{VOA}(U_{2,4},v)$ from two $v\times v$ orthogonal Latin squares. That is, for each row of $\mathbf{T}_{\rm Eu}$, the first two entries are the row and column indices of the two squares, and the third and forth entries are the symbols of the first square and the second square, respectively.
    Let $\bm{\alpha}=(1,1,1,1,2)$. Let $A_{i}=\{i-1\},i=1,2,3,4$ and $A_{5}=\{4,5\}$. For any partition $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{t})$ of $6$ other than $(1,1,1,1,1,1)$, it is coarser than $\bm{\alpha}$. Let $\{B_{i},i=1,2,\ldots,t\}$ be a partition of $N_{6}$ with $|B_{i}|=\beta_{i}$ such that each $B_{i}$ is a union of some $A_{i}$s. Let $\mathbf{T}^{\prime}_{\rm Eu}$ be a $36\times 4$ array such that $\mathbf{T}^{\prime}_{\rm Eu}(i)=\mathbf{T}_{\rm Eu}(i)$, $i=1,2,3$ and each entry in $\mathbf{T}^{\prime}_{\rm Eu}(4)$ is $j$ if the corresponding entry in $\mathbf{T}_{\rm Eu}(4)$ is in $B_{j}$. Let $\mathbf{X}=(X_{i},i\in N_{4})$ distributed on the rows of $\mathbf{T}^{\prime}_{\rm Eu}$. The entropy function $\mathbf{h}$ of $\mathbf{X}$ satisfies that $a+b=\log 6$ and $a=H(\bm{\beta})$.
The theorem is proved. ∎
Remark A frequency square induced by partition $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{t})$ of integer $v$ is a $v\times v$ square $S$ with symbols $k\in\mathbb{I}_{t}$, each appearing in each row and each column of $S$ $\alpha_{k}$ times. Note that when $\bm{\alpha}$ is the all-$1$ partition of $v$, $S$ reduces to a Latin square. A Latin square $S_{1}$ of order $v$ (and of the first kind) and a frequency square $S_{2}$ induced by partition $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{t})$ of integer $v$ are called orthogonal if each pair $(s_{1},s_{2})\in\mathbb{I}_{v}\times\mathbb{I}_{t}$ appears $\alpha_{i}$, where $s_{2}$ is the symbol that appears $\alpha_{i}$ times in each row and column of $S_{2}$. By characterizing entropy functions on the face $(U_{2,4},U^{123}_{2,3})$, this theorem also studies arrays corresponding to orthogonal two squares with one Latin square $S_{1}$ and one frequency square $S_{2}$. When $\bm{\alpha}$ is the all-$1$ partition of $v$, it reduces to orthogonal of two Latin squares and then corresponding entropy functions are on the extreme ray containing $U_{2,4}$. On the other hand, when $\bm{\alpha}$ is the trivial partition of a single $v$, it reduces to orthogonal squares of one Latin square of the first kind and the other of the zeroth kind, and entropy functions are in the extreme ray containing $U^{123}_{2,3}$. It is interesting that although $\mathrm{VOA}(U_{2,4},6)$ does not exist, arrays “between” $\mathrm{VOA}(U^{123}_{2,3},6)$ and $\mathrm{VOA}(U_{2,4},6)$ exist.
Theorem III-B.3.
For $F=(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log v$ for some positive $v$ and $a=\frac{1}{v}\sum_{i=0}^{v-1}H(\bm{\alpha_{i}})$, where $\bm{\alpha_{i}}\in\mathcal{P}(v),i\in\mathbb{I}_{v}$. Here, $\mathcal{P}(v)$ denotes the family of all partitions of $v$.
[…]
Figure 7: The face $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\quad i\in N_{4},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in\{2,3,4\},$
$H(X_{12})+H(X_{1i})$ $=H(X_{1})+H(X_{12i}),\quad i\in\{3,4\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad|K|=2,K\neq\{1,2\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.2.34)
$=p(x_{1},x_{2},x_{4})$ (III.2.35)
$=p(x_{1},x_{3},x_{4})$ (III.2.36)
$=p(x_{2},x_{3},x_{4}),$ (III.2.37)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.2.38)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.2.39)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.2.40)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (III.2.41)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (III.2.42)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.2.43)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (III.2.44)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.2.45)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.2.46)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.2.47)
Restricting $\mathbf{h}=(a,b)$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=(a+b)\mathbf{r}^{\prime}$, where $\mathbf{r}^{\prime}$ is the rank function of $U_{2,3}$ on $\{2,3,4\}$. Thus the characterizing random vector $(X_{2},X_{3},X_{4})$ of $\mathbf{h}^{\prime}$ is uniformly distributed on the rows of a $\mathrm{VOA}(U_{2,3},v)$ for a positive integer $v$, and so $a+b$ can only take the value of $\log v$. Now let $\mathbf{T}$ be a $v^{2}\times 4$ array with $\mathbf{T}(2,3,4)$ a $\mathrm{VOA}(U_{2,3},v)$. By (III.2.37), $X_{1}$ is a function of $(X_{2},X_{3},X_{4})$, which implies that $(X_{i},i\in N_{4})$ must be distributed on the rows of a such constructed $\mathbf{T}$.
According to (III.2.36), canceling $p(x_{1},x_{3},x_{4})$ on the left side and $p(x_{1},x_{2},x_{3},x_{4})$ on the right side of (III.2.43), we have
$p(x_{1},x_{2},x_{3})=p(x_{1},x_{3}).$ (III.2.48)
Then canceling $p(x_{1},x_{3})$ and $p(x_{1},x_{2},x_{3})$ in (III.2.41), we obtain
$p(x_{1},x_{2})=p(x_{1}),$ (III.2.49)
which implies $X_{2}$ is a function of $X_{1}$. Then for each $j\in\mathcal{X}_{1}$, there exists a unique $i\in\mathcal{X}_{2}=\mathbb{I}_{v}$ such that $(j,i)$ forms a row in $\mathbf{T}(1,2)$. Let $\beta_{i,j}$ denote the times of the row $(j,i)$ that occurs in $\mathbf{T}(1,2)$. Note that each $i\in\mathbb{I}_{v}$ occurs in $\mathbf{T}(2)$ exactly $v$ times by the definition of a $\mathrm{VOA}(U_{2,3},k)$. So for each $i\in\mathbb{I}_{v}$, $\bm{\beta_{i}}=(\beta_{i,j}>0,j\in\mathcal{X}_{1})$ forms a partition of $v$. We assume that there exist $t_{i}$ different $j$ such that $\beta_{i,j}>0$ for $i\in\mathbb{I}_{v}$. Then, $\bm{\beta_{i}}$ can be written as $\bm{\alpha_{i}}=(\alpha_{i,0},\dots,\alpha_{i,t_{i}-1})$. Then
$H(X_{1})=H($ $\frac{\alpha_{0,0}}{v^{2}},\frac{\alpha_{0,1}}{v^{2}},\dots,\frac{\alpha_{0,t_{1}-1}}{v^{2}},$
$\frac{\alpha_{1,0}}{v^{2}},\frac{\alpha_{1,1}}{v^{2}},\dots,\frac{\alpha_{1,t_{2}-1}}{v^{2}},$
$\frac{\alpha_{v-1,0}}{v^{2}},\frac{\alpha_{v-1,1}}{v^{2}},\dots,\frac{\alpha_{v-1,t_{v}-1}}{v^{2}})$
$=-\sum_{i=0}^{v-1}$ $\sum_{j=0}^{t_{i}-1}\frac{\alpha_{i,j}}{v^{2}}\log\frac{\alpha_{i,j}}{v^{2}}.$ (III.2.50)
As $\mathbf{h}\in F$, $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b,$ (III.2.51)
$H(X_{2})$ $=H(X_{3})=H(X_{4})=a+b.$ (III.2.52)
Since $a+b=\log v$, we have
$a$ $=H(X_{1})-(a+b)$ (III.2.53)
$=-\sum_{i=0}^{v-1}\sum_{j=0}^{t_{i}-1}\frac{\alpha_{i,j}}{v^{2}}\log\frac{\alpha_{i,j}}{v^{2}}-\log v$ (III.2.54)
$=-\sum_{i=0}^{v-1}\sum_{j=0}^{t_{i}-1}\frac{\alpha_{i,j}}{v^{2}}\log\frac{\alpha_{i,j}}{v}=\frac{1}{v}\sum_{i=0}^{v-1}H(\bm{\alpha_{i}}).$ (III.2.55)
Now we prove the “if” part. For any $\mathbf{h}=(a,b)$ satisfying $a+b=\log v$ and $a=\frac{1}{v}\sum_{i=0}^{v-1}H(\bm{\alpha_{i}})$, where $\bm{\alpha_{i}}=(\alpha_{i,0},\dots,\alpha_{i,t_{i}-1})\in\mathcal{P}(v),i\in\mathbb{I}_{v}$, let $\mathbf{T}$ be a $v^{2}\times 4$ array such that $\mathbf{T}(2,3,4)$ is a $\mathrm{VOA}(U_{2,3},v)$. In the rows where $i$ occurs in the $\mathbf{T}(2)$, let the entry $(\sum_{m=0}^{i-1}t_{m})+j$ occur $\alpha_{i,j}$ times in $\mathbf{T}(1)$ for $i\in\mathbb{I}_{v}$ and $j\in\mathbb{I}_{t_{i}}$ where we set $t_{-1}=0$. Then let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}.$ It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$. The proof is accomplished. ∎
Remark Similar to the first two faces in this Subsection, the characterization of $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$ can also be considered as two squares of order $v$, where $S_{2}$ is a Latin square and $S_{1}$ is a square whose symbols can be obtained from splitting symbols in $S_{2}$, that is, each symbol $i$ splits into $t_{i}$ symbols with each occurring $\alpha_{i,j}$ times according to the partition $\bm{\alpha}_{i}$.
III-C Entropy functions on faces involving Latin square decompositions
In this subsection, we introduce three types of decompositions of a $\mathrm{VOA}(U_{2,3},v)$, which characterize three $2$-dimensional faces, respectively. As discussed in Subsection II-C, each $\mathrm{VOA}(U_{2,3},v)$ corresponds to a Latin square, the three types of $\mathrm{VOA}(U_{2,3},v)$ decompositions correspond to three types of Latin square decomposition.
Definition 2.
Given $A,B\subseteq\mathbb{I}_{v}$ and a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$, an $|A||B|\times 3$ subarray $\mathbf{T}^{\prime}$ of $\mathbf{T}$ is called induced by $A$ and $B$ if rows in $\mathbf{T}^{\prime}(1,2)$ are exactly those pairs in $A\times B$.
Definition 3.
Given $A,B\subseteq\mathbb{I}_{v}$ with $|A||B|=v$ and a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$,
  • a subarray $\mathbf{T}^{\prime}$ of $\mathbf{T}$ induced by $A$ and $B$ is called a unit subarray of $\mathbf{T}$ if each $e\in\mathbb{I}_{v}$ occurs exactly once in $\mathbf{T}^{\prime}(3)$.
  • $\{\mathbf{T}_{i},i\in\mathbb{I}_{v}\}$ is called an uniform decomposition of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$ if
    • each $\mathbf{T}_{i}$ induced by $A_{i}$ and $B_{i}$ is a unit subarray of $\mathbf{T}$ and
    • $\biguplus\limits_{i\in\mathbb{I}_{v}}A_{i}\times B_{i}=\mathbb{I}^{2}_{v}$.
Example 4.
Here is an example of uniform decomposition of a $\mathrm{VOA}(U_{2,3},4)$ $\mathbf{T}$.
$\mathbf{T}\ =\ \begin{matrix}0&0&0\\ 0&1&3\\ 0&2&1\\ 0&3&2\\ 1&0&1\\ 1&1&2\\ 1&2&0\\ 1&3&3\\ 2&0&2\\ 2&1&0\\ 2&2&3\\ 2&3&1\\ 3&0&3\\ 3&1&1\\ 3&2&2\\ 3&3&0\\ \end{matrix}\qquad\mathbf{T}_{0}\ =\ \begin{matrix}0&0&0\\ 0&1&3\\ 0&2&1\\ 0&3&2\end{matrix}\qquad\mathbf{T}_{1}\ =\ \begin{matrix}1&0&1\\ 1&1&2\\ 1&2&0\\ 1&3&3\\ \end{matrix}\qquad\mathbf{T}_{2}\ =\ \begin{matrix}2&0&2\\ 2&1&0\\ 3&0&3\\ 3&1&1\end{matrix}\qquad\mathbf{T}_{3}\ =\ \begin{matrix}2&2&3\\ 2&3&1\\ 3&2&2\\ 3&3&0\end{matrix}$
Note that in this exmaple, $\mathbf{T}_{i}$ is induced by $A_{i}$ and $B_{i}$, where $i\in\mathbb{I}_{4}$ and
  • $A_{0}=\{0\}$, $B_{0}=\mathbb{I}_{4}$,
  • $A_{1}=\{1\}$, $B_{1}=\mathbb{I}_{4}$,
  • $A_{2}=\{2,3\}$, $B_{2}=\{0,1\}$ and
  • $A_{3}=\{2,3\}$, $B_{3}=\{2,3\}$.
Theorem III-C.1.
For $F=(\mathcal{W}_{2}^{12},\mathcal{W}_{2}^{13})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if there exists a uniform decomposition $\{\mathbf{T}_{0},\dots,\mathbf{T}_{v-1}\}$ of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$ such that
$a=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|B_{i}|\text{ and }b=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|A_{i}|,$
where the subarray $\mathbf{T}_{i}$ of $\mathbf{T}$ are induced by $A_{i}$ and $B_{i}$ for $i\in\mathbb{I}_{v}$.
[…]
Figure 8: The face $(\mathcal{W}_{2}^{12},\mathcal{W}_{2}^{13})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),$
$\qquad\qquad\{i,j\}\neq\{1,2\},\{1,3\}$
$H(X_{12})+H(X_{13})$ $=H(X_{1})+H(X_{123}),$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad|K|=2,K\neq\{1,2\},\{1,3\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.3.1)
$=p(x_{1},x_{2},x_{4})$ (III.3.2)
$=p(x_{1},x_{3},x_{4})$ (III.3.3)
$=p(x_{2},x_{3},x_{4})$ (III.3.4)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.3.5)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.3.6)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.3.7)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.3.8)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (III.3.9)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}),$ (III.3.10)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.3.11)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234})$ (III.3.12)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.3.13)
Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=(a+b)\mathbf{r}^{\prime}$, where $\mathbf{r}^{\prime}$ is the rank function of $U_{2,3}$ on $\{2,3,4\}$. Thus the characterizing random vector $(X_{2},X_{3},X_{4})$ of $\mathbf{h}^{\prime}$ is uniformly distributed on the rows of a $\mathrm{VOA}(U_{2,3},v)$ for a positive integer $v$, and so $a+b$ can only take the value of $\log v$.
By (III.3.1)-(III.3.4), (III.3.10) and (III.3.12), we obtain $p(x_{1},x_{4})=p(x_{2},x_{4})$. Then with (III.3.5) and (III.3.7), we have $p(x_{1})p(x_{4})=p(x_{2})p(x_{4})$. Therefore, we obtain $p(x_{1})=p(x_{2})=\frac{1}{v}$, which implies that $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$ and $H(X_{1})=\log v$.
Since $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$, $(X_{i},i\in N_{4}$) must be uniformly distributed on the rows of a $v^{2}\times 4$ array $\mathbf{T}$ such that $\mathbf{T}(2,3,4)$ is a $\mathrm{VOA}(U_{2,3},v)$, and the first entry of each row in $\mathbf{T}$ is uniquely determined by the remaining three entries. Assume $\mathcal{X}_{1}=\mathbb{I}_{v}$. Let $A_{i}$ and $B_{i}$ denote the set of all $j_{1}$ such that $(i,j_{1})$ appears on the row of $\mathbf{T}(1,2)$, and the set of $j_{2}$ such that $(i,j_{2})$ appears on the row of $\mathbf{T}(1,3)$ for $i\in\mathbb{I}_{v}$, respectively. For any $j_{1}\in A_{i}$ and $j_{2}\in B_{i}$, we have
$p_{X_{1},X_{2}}(i,j_{1})>0\text{ and }p_{X_{1},X_{3}}(i,j_{2})>0.$ (III.3.14)
Together with (III.3.9), we obtain
$p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})>0.$ (III.3.15)
Note that $(j_{1},j_{2})$ occurs exactly once on the rows of $\mathbf{T}(2,3)$ due to the definition of $\mathrm{VOA}(U_{2,3},v)$, $(i,j_{1},j_{2})$ will appear exactly once on the rows of $\mathbf{T}(1,2,3)$, and so
$p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})=\frac{1}{v^{2}}.$ (III.3.16)
The probability mass of $(i,j_{1})$, $j_{i}\in A_{i},i\in\mathbb{I}_{v}$
$p_{X_{1},X_{2}}(i,j_{1})=\sum_{j_{2}\in B_{i}}p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})=\frac{|B_{i}|}{v^{2}}.$ (III.3.17)
Similarly, the probability mass of $(i,j_{2})$, $j_{2}\in B_{i},i\in\mathbb{I}_{v}$
$p_{X_{1},X_{3}}(i,j_{2})=\sum_{j_{1}\in A_{i}}p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})=\frac{|A_{i}|}{v^{2}}.$ (III.3.18)
Since $p(x_{1})=\frac{1}{v}$, together with (III.3.9) and (III.3.16)-(III.3.18), we obtain
$|A_{i}|\times|B_{i}|=v,$ (III.3.19)
which implies that $i$ will occur $v$ times in $\mathbf{T}(1)$. Note $X_{1}$ is independent of $X_{4}$ and $|\mathcal{X}_{4}|=v$, $(i,j_{3})$ for $j_{3}\in N_{4}$ appears exactly once on the rows of $\mathbf{T}(1,4)$. Thus $\mathbf{T}$ can be decomposed into $v$ arrays based on the entry $i$ occuring on the rows of $\mathbf{T}(1)$, that is, the entries on the rows that $i\in\mathbf{T}(1)$ appears of $\mathbf{T}$ forms a $v\times 4$ array $\mathbf{T}_{i}$ for $i\in\mathbb{I}_{v}$. We can check that $\{\mathbf{T}_{i}(2,3,4),i\in\mathbb{I}_{v}\}$ is a uniform decomposition of $\mathbf{T}(2,3,4)$, and the entries of $\mathbf{T}_{i}(2)$$\mathbf{T}_{i}(3)$ and $\mathbf{T}_{i}(4)$ are from $A_{i}$, $B_{i}$ and $\mathbb{I}_{v}$, respectively. The entropy of $(X_{1},X_{2})$
$H(X_{1},X_{2})$ $=H(\underbrace{\dfrac{|B_{0}|}{v^{2}},\dots,\dfrac{|B_{0}|}{v^{2}}}_{|A_{0}|},\dots,\underbrace{\dfrac{|B_{v-1}|}{v^{2}},\dots,\dfrac{|B_{v-1}|}{v^{2}}}_{|A_{v-1}|})$ (III.3.20)
$=2\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|B_{i}|.$ (III.3.21)
As $\mathbf{h}\in F$, $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=a+b$ (III.3.22)
$H(X_{1},X_{2})$ $=2a+b$ (III.3.23)
Note that $a+b=\log v$, we have
$a$ $=2\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|B_{i}|-\log k$ (III.3.24)
$=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|B_{i}|.$ (III.3.25)
By the same argument,
$b=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|A_{i}|.$ (III.3.26)
As for the “if” part, let $(X_{2},X_{3},X_{4})$ be uniformly distributed on the rows of the $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$. Let $X_{1}$ be $i$ if $(x_{2},x_{3},x_{4})$ appears on the rows of $\mathbf{T}_{i}$. Then $(X_{i},i\in N_{4})$ characterizes $(a,b)$. The proof has been completed. ∎
Remark Theorem III-C.1 establishes a correspondence between the $2$-dim face characterization problem and the uniform decomposition problem of a $\mathrm{VOA}(U_{2,3},v)$. When $v$ is prime, $\mathrm{VOA}(U_{2,3},v)$ can be decomposed into $v$ uniform subarrays where either $|A_{i}|=1$ and $|B_{i}|=v$ for $i\in\mathbb{I}_{v}$, or $|A_{i}|=v$ and $|B_{i}|=1$ for $i\in\mathbb{I}_{v}$. These correspond to the polymatroids $(0,\log v)$ and $(\log v,0)$, respectively. While for a composite $v$, the uniform decomposition of a $\mathrm{VOA}(U_{2,3},v)$ will be more complicated. In Example 4, the uniform decomposition corresponds to the entropy function $(0.5,1.5)$ on the face $F$.
Definition 4.
Given $A,B\subseteq\mathbb{I}_{v}$ with $|A|=|B|=v^{\prime}\leq v$ and a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$,
  • a subarray $\mathbf{T}^{\prime}$ of $\mathbf{T}$ induced by $A$ and $B$ is called a suborder $\mathrm{VOA}$ of $\mathbf{T}$ if $\mathbf{T}^{\prime}$ is a $\mathrm{VOA}(U_{2,3},v^{\prime})$.
  • $\{\mathbf{T}_{i},i\in\mathbb{I}_{t}\}$ is called a suborder decomposition of $\mathbf{T}$ if
    • each $\mathbf{T}_{i}$ induced by $A_{i}$ and $B_{i}$ is a suborder $\mathrm{VOA}$ of $\mathbf{T}$ and
    • $\biguplus\limits_{i\in\mathbb{I}_{t}}A_{i}\times B_{i}=\mathbb{I}^{2}_{v}$.
Example 5.
Given a $\mathrm{VOA}(U_{2,3},4)$ $\mathbf{T}$ in the following, it can be seen that $\{\mathbf{T}_{0},\mathbf{T}_{1},\ldots,\mathbf{T}_{6}\}$ forms a suborder $\mathrm{VOA}$ decomposition of $\mathbf{T}$.
[…]
where
  • $A_{0}=\{0,1\}$, $B_{0}=\{0,2\}$,
  • $A_{1}=\{0,1\}$, $B_{1}=\{1,3\}$,
  • $A_{2}=\{2,3\}$, $B_{1}=\{0,2\}$,
  • $A_{3}=\{2\}$, $B_{3}=\{1\}$,
  • $A_{4}=\{2\}$, $B_{4}=\{3\}$,
  • $A_{5}=\{3\}$, $B_{3}=\{1\}$ and
  • $A_{6}=\{3\}$, $B_{4}=\{3\}$.
Theorem III-C.2.
For $F=(\hat{U}_{2,5}^{1},U^{234}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log v$ for some positive $v$ and there exists a suborder decomposition $\{\mathbf{T}_{0},\mathbf{T}_{1},\ldots,\mathbf{T}_{t-1}\}$ of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}$ such that
$a=\frac{1}{2}H(\dfrac{|A_{i}|^{2}}{k^{2}}:i\in\mathbb{I}_{t}),$
where subarray $\mathbf{T}_{i}$ of $\mathbf{T}$ are induced by $A_{i}$ and $B_{i}$ for $i\in\mathbb{I}_{t}$.
[…]
Figure 9: The face $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in\{2,3,4\},$
$H(X_{1i})+H(X_{1j})$ $=H(X_{1})+H(X_{1ij}),i,j\in\{2,3,4\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad|K|=2,K\subseteq\{2,3,4\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.3.27)
$=p(x_{1},x_{2},x_{4})$ (III.3.28)
$=p(x_{1},x_{3},x_{4})$ (III.3.29)
$=p(x_{2},x_{3},x_{4}),$ (III.3.30)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.3.31)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.3.32)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.3.33)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (III.3.34)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (III.3.35)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (III.3.36)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.3.37)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.3.38)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.3.39)
Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=(a+b)\mathbf{r}^{\prime}$, where $\mathbf{r}^{\prime}$ is the rank function of $U_{2,3}$ on $\{2,3,4\}$. Thus the characterizing random vector $(X_{2},X_{3},X_{4})$ of $\mathbf{h}^{\prime}$ is uniformly distributed on the rows of a $\mathrm{VOA}(U_{2,3},v)$ for a positive integer $v$, and so $a+b$ can only take the value of $\log v$. By (III.3.30), $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$, which implies that $(X_{i},i\in N_{4})$ must be uniformly distributed on the rows of a $v^{2}\times 4$ array $\mathbf{T}$ with $\mathbf{T}(2,3,4)$ a $\mathrm{VOA}(U_{2,3},v)$. Assume $\mathcal{X}_{1}=\mathbb{I}_{t}$. Let $A_{i}$ and $B_{i}$ denote the set of all $j_{1}$ such that $(i,j_{1})$ appears on the row of $\mathbf{T}(1,2)$, and the set of $j_{2}$ such that $(i,j_{2})$ appears on the row of $\mathbf{T}(1,3)$ for $i\in\mathbb{I}_{t}$, respectively. For any $j_{1}\in A_{i}$ and $j_{2}\in B_{i}$, we have
$p_{X_{1},X_{2}}(i,j_{1})>0,p_{X_{1},X_{3}}(i,j_{2})>0.$ (III.3.40)
Together with (III.3.34), we obtain
$p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})>0.$ (III.3.41)
Note that $(j_{1},j_{2})$ occurs exactly once on the rows of $\mathbf{T}(2,3)$ due to the definition of $\mathrm{VOA}(U_{2,3},v)$, $(i,j_{1},j_{2})$ will appear exactly once on the rows of $\mathbf{T}(1,2,3)$, and so
$p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})=\frac{1}{v^{2}}.$ (III.3.42)
The probability mass of $(i,j_{1})$ for $i\in\mathbb{I}_{t},j_{1}\in A_{i}$
$p_{X_{1},X_{2}}(i,j_{1})=\sum_{j_{2}\in B_{i}}p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})=\frac{|B_{i}|}{v^{2}}.$ (III.3.43)
Similarly, the probability mass of $(i,j_{2})$ for $i\in\mathbb{I}_{t},j_{2}\in B_{i}$
$p_{X_{1},X_{3}}(i,j_{2})=\sum_{j_{1}\in A_{i}}p_{X_{1},X_{2},X_{3}}(i,j_{1},j_{2})=\frac{|A_{i}|}{v^{2}}.$ (III.3.44)
By (III.3.27)-(III.3.29), $p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})=p(x_{1},x_{3},x_{4})$. Equating the left side of (III.3.34)-(III.3.36), we obatin
$p(x_{1},x_{2})p(x_{1},x_{3})=p(x_{1},x_{2})p(x_{1},x_{4})=p(x_{1},x_{3})p(x_{1},x_{4}),$ (III.3.45)
which implies
$p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (III.3.46)
Together with (III.3.43) and (III.3.44), we obtain
$|A_{i}|=|B_{i}|.$ (III.3.47)
Assume $E_{i}$ denotes the set of $j_{3}$ satisfying $(i,j_{3})$ appears on the row of $\mathbf{T}(1,4)$ for $i\in\mathbb{I}_{t}$. By symmetry, we conclude
$|A_{i}|=|B_{i}|=|E_{i}|.$ (III.3.48)
Note that $\mathbf{T}(2,3,4)$ is a $\mathrm{VOA}(U_{2,3},v)$, the subarray $\mathbf{T}_{i}$ of $\mathbf{T}(2,3,4)$ induced by $A_{i}$ and $B_{i}$ is a $\mathrm{VOA}(U_{2,3},|A_{i}|)$. Therefore, each $i\in N_{t}$ determines a suborder $\mathrm{VOA}$ $\mathbf{T}_{i}$ of $\mathbf{T}(2,3,4)$, which implies $\{\mathbf{T}_{i},i\in\mathbb{I}_{t}\}$ is a suborder $\mathrm{VOA}$ decomposition of $\mathbf{T}(2,3,4)$. The probalities mass of $i\in\mathbb{I}_{t}$
$p_{X_{1}}(i)=\sum_{j_{1}\in A_{i}}p_{X_{1},X_{2}}(i,j_{1})$ (III.3.49)
$=\frac{|A_{i}||B_{i}|}{v^{2}}=\frac{|A_{i}|^{2}}{v^{2}}.$ (III.3.50)
Thus the entropy of $X_{1}$ is equal to
$H(X_{1})=H(\dfrac{|A_{i}|^{2}}{v^{2}}:i\in\mathbb{I}_{t})$ (III.3.51)
As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, restricting $\mathbf{h}$ on $\{1\}$, we have
$H(X_{1})$ $=2a.$ (III.3.52)
which implies
$a=\frac{1}{2}H(\dfrac{|A_{i}|^{2}}{v^{2}}:i\in\mathbb{I}_{t}).$ (III.3.53)
For the “if” part, let $(X_{2},X_{3},X_{4})$ be uniformly distributed on the rows of a $\mathrm{VOA}(U_{2,3},v)$ $\mathbf{T}^{\prime}$. Let $\{\mathbf{T}_{0},\ldots,\mathbf{T}_{t-1}\}$ be a suborder $\mathrm{VOA}$ decomposition of $\mathbf{T}^{\prime}$. Let $X_{1}$ be $i$ if $(x_{2},x_{3},x_{4})$ appear on the rows of $\mathbf{T}_{i}$. Then the entropy function $\mathbf{h}$ of $(X_{i},i\in N_{4})$ is in $F$. The proof has been completed. ∎
Remark Theorem III-C.2 establishes a correspondence between the face $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$ characterization and suborder decomposition problem. It is obvious that $\mathrm{VOA}(U_{2,3},v)$ is inherently a suborder $\mathrm{VOA}$ of itself. On the other hand, any $\mathrm{VOA}(U_{2,3},v)$ can be decomposed into $v^{2}$ suborder VOA $\mathrm{VOA}(U_{2,3},1)$. These two cases correspond to the polymatroids $(0,\log v)$ and $(\log v,0)$, respectively. However, listing all the $\mathrm{VOA}$ decompositions of a $\mathrm{VOA}(U_{2,3},v)$ can be challenging.
Definition 5.
For a $v^{2}\times 4$ array $\mathbf{T}$, if
  • $\mathbf{T}(2,3,4)$ is a $\mathrm{VOA}(U_{2,3},v)$,
  • entries in $\mathbf{T}(1)$ is from $\mathbb{I}_{t}$ with $v\leq t\leq v^{2}$, and
  • for each $i=2,3,4$, each row in $\mathbf{T}(1,i)$ occurs exactly once,
we call $\mathbf{T}$ a $\{1\}$-partial $\mathrm{VOA}(U_{2,4})$.
Example 6.
Let
$\mathbf{T}=\begin{matrix}0&0&0&0\\ 2&0&1&1\\ 1&0&2&2\\ 3&1&0&2\\ 1&1&1&0\\ 0&1&2&1\\ 1&2&0&1\\ 4&2&1&2\\ 2&2&2&0\end{matrix}$
It can be seen that $\mathbf{T}$ is a $\{1\}$-partial $\mathrm{VOA}(U_{2,4})$, and the entries of $\mathbf{T}(1)$ is from $\mathbb{I}_{5}$.
Theorem III-C.3.
For $F=(\hat{U}_{2,5}^{1},U_{2,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log v$ and there exists a $\{1\}$-partial $\mathrm{VOA}(U_{2,4},v)$ such that
$a=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}})-\log v,$
where $\alpha_{i}$ denotes the times of the entry $i\in\mathbb{I}_{t}$ that occurs in $\mathbf{T}(1)$.
[…]
Figure 10: The face $(\hat{U}_{2,5}^{1},U_{2,4})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in\{2,3,4\}$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),|K|=2.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalties imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.3.54)
$=p(x_{1},x_{2},x_{4})$ (III.3.55)
$=p(x_{1},x_{3},x_{4})$ (III.3.56)
$=p(x_{2},x_{3},x_{4})$ (III.3.57)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.3.58)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.3.59)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.3.60)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.3.61)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.3.62)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (III.3.63)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.3.64)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.3.65)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.3.66)
By (III.3.54), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.3.61), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.3.67)
Together with (III.3.55), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (III.3.68)
By the same argument, we have
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{3})=p(x_{1},x_{4})$ (III.3.69)
Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=(a+b)\mathbf{r}^{\prime}$, where $\mathbf{r}^{\prime}$ is the rank function of $U_{2,3}$ on $\{2,3,4\}$. Thus the characterizing random vector $(X_{2},X_{3},X_{4})$ of $\mathbf{h}^{\prime}$ is uniformly distributed on the rows of a $\mathrm{VOA}(U_{2,3},v)$ for a positive integer $k$, and so $a+b$ can only take the value of $\log v$.
By (III.3.57), $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$, which implies that $(X_{i},i\in N_{4})$ must be uniformly distributed on the rows of a $v^{2}\times 4$ array $\mathbf{T}$ with $\mathbf{T}(2,3,4)$ a $\mathrm{VOA}(U_{2,3},v)$. Note that $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})$ by (III.3.68), each row of $\mathbf{T}(1,2)$ occurs exactly once in $\mathbf{T}(1,2)$. Similarly, by (III.3.69), each row of $\mathbf{T}(A)$ occurs exactly once in $\mathbf{T}(A)$ for $A=\{1,3\}$ and $M=\{1,4\}$. Hence, $\mathbf{T}$ is a $\{1\}$-partial $\mathrm{VOA}(U_{2,4})$. Recall that $(X_{i},i\in N_{4})$ is uniformly distributed on the rows of $\mathbf{T}$, the probability of each row $(x_{1},x_{2},x_{3},x_{4})$ of $\mathbf{T}$ is
$p(x_{1},x_{2},x_{3},x_{4})=\frac{1}{v^{2}}.$ (III.3.70)
Then
$H(X_{1})=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}}),$ (III.3.71)
where $\alpha_{i}$ denotes the times of $i\in\mathbb{I}_{t}$ that occurs in $\mathbf{T}(1)$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b.$ (III.3.72)
$H(X_{2})$ $=H(X_{3})=H(X_{4})=a+b$ (III.3.73)
Note that $a+b=\log v$, we obtain that
$a=H(\dfrac{\alpha_{0}}{v^{2}},\dfrac{\alpha_{1}}{v^{2}},\dots,\dfrac{\alpha_{t-1}}{v^{2}})-\log v.$ (III.3.74)
To prove the “if” part of the theorem, let $\mathbf{T}$ be a $\{1\}$-partial $\mathrm{VOA}(U_{2,4})$. Let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}$. Then $(X_{i},i\in N_{4})$ characterizes $(a,b)$. The proof is accomplished. ∎
III-D Entropy functions on the face $(\hat{U}_{3,5}^{4},U_{2,4})$
In this subsection, we characterize entropy functions on the face $(\hat{U}_{3,5}^{4},U_{2,4})$, which is a face with one extreme ray containing a rank 3 integer polymatroid and another containing a rank 2 matroid.
Theorem III-D.1.
For $F=(\hat{U}_{3,5}^{4},U_{2,4})$, $\mathbf{h}=(a,b)\in F$ is
  • entropic if
    • $\ast$
      $a+b=\log v$ for integer $v\neq 2,6$;
    • $\ast$
      $(a,b)=(\log 2,0)$; or
    • $\ast$
      $a+b=\log 6,a\geq\log 2$; and
  • non-entropic if
    • $\ast$
      $a+b\neq\log v$ for some integer $v>0$;
    • $\ast$
      $a+b=\log 2$, $a<\log 2$; or
    • $\ast$
      $(a,b)=(0,\log 6)$.
[…]
Figure 11: The face $(\hat{U}_{3,5}^{4},U_{2,4})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i,j\in N_{4},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\quad|K|=2,\{4\}\subseteq K\subseteq N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.4.1)
$=p(x_{1},x_{2},x_{4})$ (III.4.2)
$=p(x_{1},x_{3},x_{4})$ (III.4.3)
$=p(x_{2},x_{3},x_{4}),$ (III.4.4)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (III.4.5)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (III.4.6)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (III.4.7)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.4.8)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.4.9)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.4.10)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}),$ (III.4.11)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.4.12)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.4.13)
By (III.4.1), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1234})$ on either side of (III.4.11), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{4}).$ (III.4.14)
Together with (III.4.2), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{4}).$ (III.4.15)
By the same argument,
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})=p(x_{3},x_{4}).$ (III.4.16)
In light of (III.4.7), (III.4.9) and (III.4.10), together with (III.4.15) and (III.4.16),
$p(x_{1})p(x_{4})=p(x_{2})p(x_{4})=p(x_{3})p(x_{4}),$ (III.4.17)
which impies $p(x_{1})=p(x_{2})=p(x_{3})$. Since $X_{1}$, $X_{2}$, and $X_{3}$ are pairwise independent by (III.4.5), (III.4.6) and (III.4.8), by Lemma 1, we obtain that $X_{i}$ are uniformly distributed on $\mathcal{X}_{i}$ for $i=1,2,3$, and so $H(X_{1})=H(X_{2})=H(X_{3})=\log v$ where $v=|\mathcal{X}_{1}|=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$. As $\mathbf{h}\in F$, $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=H(X_{2})=H(X_{3})=a+b,$ (III.4.18)
$H(X_{4})$ $=2a+b,$ (III.4.19)
$H(X_{2},X_{4})$ $=3a+2b,$ (III.4.20)
$H(X_{1},X_{2},X_{3})$ $=3a+2b.$ (III.4.21)
Thus $a+b$ must take the value of $\log v$.
Now we give all the construction of $(a,b)$ on the region $a+b=\log v$ for $v\neq 2,6$. Let $p_{i}>0$, $i\in\mathbb{I}_{v}$ such that $H(p_{0},p_{1},\dots,p_{v-1})=a$. Let $\mathbf{T}_{0}$ be a $\mathrm{VOA}(U_{2,4},v)$. Let $\mathbf{T}_{i}$, $i=1,\dots,v-1$ be a $v^{2}\times 4$ array such that
$\mathbf{T}_{i}(1,2)$ $=\mathbf{T}_{0}(1,2),$ (III.4.22)
$\mathbf{T}_{i}(3;j)$ $=\mathbf{T}_{0}(3;j)+i\mod v\text{ for }j\in N_{k^{2}}.$ (III.4.23)
$\mathbf{T}_{i}(4;j)$ $=\mathbf{T}_{0}(4;j)+vi\text{ for }j\in N_{k^{2}},$ (III.4.24)
It can be seen that each of such constructed $\mathbf{T}_{i}$ is a $\mathrm{VOA}(U_{2,4},v)$. Let $(X_{i},i\in N_{4})$ be distributed on the rows of $\mathbf{T}\triangleq\left[\begin{matrix}\mathbf{T}_{0},\\ \dots\\ \mathbf{T}_{v-1}\end{matrix}\right]$ such that the probability mass of each row of $\mathbf{T}_{i}$ is $\frac{p_{i}}{v^{2}}$. Now we show the entropy function of $(X_{i},i\in N_{4})$ is $(a,b)$. By (III.4.24), $|\mathcal{X}_{4}|=v^{2}$ and each entry $j\in\mathcal{X}_{4}$ occurs only in one of the arrays $\mathbf{T}_{0}(4)$, $\dots$, $\mathbf{T}_{v-1}(4)$, which implies that for any subset $A$ such that $\{4\}\subseteq A\subseteq N_{4}$, each $x_{A}\in\mathcal{X}_{A}$ appears $\Delta_{A}$ times, where
$\Delta_{A}:=v^{\mathbf{r}_{U_{2,4}}(N_{4})-\mathbf{r}_{U_{2,4}}(A)},$
and $\mathbf{r}_{U_{2,4}}$ is the rank function of $U_{2,4}$. So
$H(X_{A})$ $=H\Big(\underbrace{\frac{p_{0}\Delta_{A}}{v^{2}},\dots,\frac{p_{0}\Delta_{A}}{v^{2}}}_{v^{\mathbf{r}_{U_{2,4}}(A)}},\dots,$
$\qquad\underbrace{\frac{p_{v-1}\Delta_{A}}{v^{2}},\dots,\frac{p_{v-1}\Delta_{A}}{v^{2}}}_{v^{\mathbf{r}_{U_{2,4}}(A)}}\Big)$ (III.4.25)
$=H(p_{0},p_{1},\dots,p_{v-1})+\log v\cdot\mathbf{r}_{U_{2,4}}(A)$ (III.4.26)
$=a+(a+b)\cdot\mathbf{r}_{U_{2,4}}(A)$ (III.4.27)
$=a\cdot(1+\mathbf{r}_{U_{2,4}}(A))+b\cdot\mathbf{r}_{U_{2,4}}(A).$ (III.4.28)
Then it is clear that $H(X_{1})=H(X_{2})=\log{v}$ and $H(X_{1},X_{2})=2\log{v}$. Since each $j\in\mathcal{X}_{3}$ appears exactly $v$ times in each $\mathbf{T}_{i}$, $i=0,\dots v-1$, we obtain
$H(X_{3})$ $=H(\frac{v(p_{0}+p_{1}+\dots,+p_{v-1})}{v^{2}},\dots,$
$\frac{v(p_{0}.+p_{1}+\dots,+p_{v-1})}{v^{2}})$ (III.4.29)
$=\log{v}.$ (III.4.30)
Each $(i,j)\in\mathcal{X}_{1}\times\mathcal{X}_{3}$ or $\mathcal{X}_{2}\times\mathcal{X}_{3}$ appears exactly once in $\mathbf{T}_{0}$, $\dots$, and $\mathbf{T}_{v-1}$, which implies
$H(X_{1},X_{3})$ $=H(X_{2},X_{3})=H(\frac{p_{0}+p_{1}+\dots,+p_{v-1}}{v^{2}},\dots,$
$\frac{p_{0}+p_{1}+\dots,+p_{v-1}}{v^{2}})$ (III.4.31)
$=2\log{v}.$ (III.4.32)
It remains to verify $H(X_{1},X_{2},X_{3})$. Each $(i,j,k)\in\mathcal{X}_{123}$ appears once in the rows of $\mathbf{T}(1,2,3)$, and so
$H(X_{1},X_{2},X_{3})$ $=H(\underbrace{\frac{p_{0}}{k^{2}},\frac{p_{0}}{v^{2}},\dots,\frac{p_{0}}{v^{2}}}_{v^{2}},\dots,$
$\qquad\underbrace{\frac{p_{v-1}}{v^{2}},\frac{p_{v-1}}{v^{2}},\dots,\frac{p_{v-1}}{v^{2}}}_{v^{2}})$ (III.4.33)
$=H(p_{0},\dots,p_{v-1})+2\log{v}$ (III.4.34)
$=a+2(a+b)=3a+2b.$ (III.4.35)
Now we show that all polymatroids with $a+b=\log 2$ are non-entropic except for $(\log 2,0)$. Assume $\mathbf{h}=(a,b)$ is entropic with $a+b=\log 2$ and $(X_{i},i\in N_{4})$ is its characterizing random vector. By the discussion above, assume without loss of generality that $X_{1},X_{2}$ and $X_{3}$ are all uniformly distributed on $\mathbb{I}_{2}$. Let $\mathbf{T}_{0}$ be the array consisting of rows being all $3$-tuples with entries in $\mathbb{I}_{2}$, that is,
$\begin{matrix}0&0&0\\ 0&0&1\\ 0&1&0\\ 0&1&1\\ 1&0&0\\ 1&0&1\\ 1&1&0\\ 1&1&1\end{matrix}$
It can be seen that $(X_{1},X_{2},X_{3})$ must be distributed on the rows of $\mathbf{T}_{0}$. Let the probatility of the $i$th-row of $\mathbf{T}_{0}$ be $p_{i}$. Since $X_{1}$, $X_{2}$, and $X_{3}$ are uniform and pairwise independent, each of $(x_{1},x_{2})$, $(x_{1},x_{3})$ and $(x_{2},x_{3})$ has probability $\frac{1}{4}$, and so
$p_{1}+p_{2}=\dfrac{1}{4},p_{3}+p_{4}=\dfrac{1}{4},p_{5}+p_{6}=\dfrac{1}{4},p_{7}+p_{8}=\dfrac{1}{4},$
$p_{1}+p_{5}=\dfrac{1}{4},p_{2}+p_{6}=\dfrac{1}{4},p_{3}+p_{7}=\dfrac{1}{4},p_{4}+p_{8}=\dfrac{1}{4},$
$p_{1}+p_{3}=\dfrac{1}{4},p_{2}+p_{4}=\dfrac{1}{4},p_{5}+p_{7}=\dfrac{1}{4},p_{6}+p_{8}=\dfrac{1}{4}.$
Solving above equations, we obtain
$p_{1}=p_{4}=p_{6}=p_{7},$ (III.4.36)
$p_{2}=p_{3}=p_{5}=p_{8}.$ (III.4.37)
Assume that either (III.4.36) or (III.4.37) vanishes, then $\mathbf{T}_{0}$ degenerates a $\mathrm{VOA}(U_{2,3},2)$ $\mathbf{T}_{1}$ and $(X_{1},X_{2},X_{3})$ is uniformly distributed on the rows of $\mathbf{T}_{1}$. Together with (III.4.18) and (III.4.21),
$a=0,b=\log 2,$ (III.4.38)
which contradicts the fact that $\mathbf{h}=(0,\log 2)$ is non-entropic. Hence, both (III.4.36) and (III.4.37) must be positive. By (III.4.1), $X_{4}$ is a function of $(X_{1},X_{2},X_{3})$, which implies that $(X_{i},i\in N_{4})$ must be distributed on a $\mathbf{T}$ such that $\mathbf{T}(1,2,3)=\mathbf{T}_{0}$. By (III.4.15) and (III.4.16), for each $x_{4}\in\mathcal{X}_{4},x_{j}\in\mathcal{X}_{j}$, $(x_{j},x_{4})$ appear exactly once on the row of $\mathbf{T}(j,4)$ for $j=1,2,3$. Additionally, $X_{4}$ is independent of $X_{i},i=1,2,3$ by (III.4.7), (III.4.9), and (III.4.10). There exists a unique $\mathbf{T}$ satisfying the above information equalities up to symmetry, and
$\mathbf{T}\ =\ \begin{matrix}0&0&0&0\\ 0&0&1&1\\ 0&1&0&2\\ 0&1&1&3\\ 1&0&0&3\\ 1&0&1&2\\ 1&1&0&1\\ 1&1&1&0\end{matrix}$
Calculating the entropy of $X_{1}$, $X_{2}$ and $(X_{1},X_{2})$, we obtain
$H(X_{2})$ $=\log 2,$ (III.4.39)
$H(X_{4})$ $=\log 4,$ (III.4.40)
$H(X_{2},X_{4})$ $=H(p_{1},p_{2},\dots,p_{8}).$ (III.4.41)
Together with (III.4.19)-(III.4.20), we have
$a=\log 2,b=0,$ (III.4.42)
$p_{1}=p_{2}=\dots=p_{8}=\dfrac{1}{8}.$ (III.4.43)
Thus the entropy function of $(X_{i},i\in N_{4})$ must be $(\log 2,0)$, which implies that all polymatroids on the region $a+b=\log 2$ are non-entropic expect for $(\log 2,0)$.
Now we show an inner bound on the entropy region on $F$ that $a+b=\log 6$ and $\log 2\leq a\leq\log 6$. Let $p_{i}>0$, $i=0,1,2$, and $p_{0}+p_{1}+p_{2}=1$ and $H(p_{0},p_{1},p_{2})=a-\log 2$. Let $\mathbf{T}^{(k)}_{0},k=1,2$ be arrays defined as follows.
$\mathbf{T}^{(1)}_{0}\ =\ \begin{matrix}0&0&0&0\\ 0&1&1&5\\ 0&2&2&3\\ 0&3&3&10\\ 0&4&4&2\\ 0&5&5&1\\ 1&0&1&1\\ 1&1&2&0\\ 1&2&5&4\\ 1&3&4&3\\ 1&4&0&5\\ 1&5&3&2\\ 2&0&2&2\\ 2&1&3&11\\ 2&2&0&1\\ 2&3&1&4\\ 2&4&5&3\\ 2&5&4&0\\ 3&0&3&3\\ 3&1&5&2\\ 3&2&4&5\\ 3&3&2&1\\ 3&4&1&0\\ 3&5&0&4\\ 4&0&4&4\\ 4&1&0&3\\ 4&2&1&2\\ 4&3&5&0\\ 4&4&3&1\\ 4&5&2&5\\ 5&0&5&5\\ 5&1&4&1\\ 5&2&3&0\\ 5&3&0&2\\ 5&4&2&4\\ 5&5&1&3\end{matrix}\qquad\qquad\mathbf{T}^{(2)}_{0}\ =\ \begin{matrix}0&0&2&6\\ 0&1&3&4\\ 0&2&0&9\\ 0&3&1&11\\ 0&4&5&8\\ 0&5&4&7\\ 1&0&3&7\\ 1&1&0&6\\ 1&2&4&10\\ 1&3&5&9\\ 1&4&2&11\\ 1&5&1&8\\ 2&0&0&8\\ 2&1&1&10\\ 2&2&2&7\\ 2&3&3&5\\ 2&4&4&9\\ 2&5&5&6\\ 3&0&1&9\\ 3&1&4&8\\ 3&2&5&11\\ 3&3&0&7\\ 3&4&3&6\\ 3&5&2&10\\ 4&0&5&10\\ 4&1&2&9\\ 4&2&3&8\\ 4&3&4&6\\ 4&4&1&7\\ 4&5&0&11\\ 5&0&4&11\\ 5&1&5&7\\ 5&2&1&6\\ 5&3&2&8\\ 5&4&0&10\\ 5&5&3&9\end{matrix}$
[…]
$\mathbf{T}^{(k)}_{i}(1,2)$ $=\mathbf{T}^{(k)}_{1}(1,2),$ (III.4.44)
$\mathbf{T}^{(k)}_{i}(3;j)$ $=\sigma_{i}(\mathbf{T}^{(k)}_{1}(3;j))\text{ for }j\in N_{36}$ (III.4.45)
$\mathbf{T}^{(k)}_{i}(4;j)$ $=\mathbf{T}^{(k)}_{1}(4;j)+12i\text{ for }j\in N_{36},$ (III.4.46)
Let
$\mathbf{T}=\left[\begin{matrix}\mathbf{T}^{(1)}_{0}\\ \mathbf{T}^{(2)}_{0}\\ \mathbf{T}^{(1)}_{1}\\ \mathbf{T}^{(2)}_{1}\\ \mathbf{T}^{(1)}_{2}\\ \mathbf{T}^{(2)}_{2}\end{matrix}\right$]
Let $(X_{i},i\in N_{4})$ be distributed on the rows of $\mathbf{T}$ such that the probability mass of each row of $\mathbf{T}^{(k)}_{i}$ is equal to $\frac{p_{i}}{72}$ for $i=0,1,2,k=1,2$.
Now we show the entropy function of $(X_{i},i\in N_{4})$ is $(a,b)$. Since entry $j+12i$, $j\in\mathbb{I}_{12}$ occurs $6$ times in $\left[\begin{matrix}\mathbf{T}^{(1)}_{i}(4)\\ \mathbf{T}^{(2)}_{i}(4)\end{matrix}\right]$ for $i=0,1,2$.
$H(X_{4})$ $=H(\underbrace{\frac{p_{0}}{12},\frac{p_{0}}{12},\dots,\frac{p_{0}}{12}}_{12},\dots,$ (III.4.47)
$\qquad\underbrace{\frac{p_{2}}{12},\frac{p_{2}}{12},\dots,\frac{p_{2}}{12}}_{12})$ (III.4.48)
$=H(p_{0},p_{1},p_{2})+\log{12}$ (III.4.49)
$=a-\log 2+\log 12$ (III.4.50)
$=2a+b.$ (III.4.51)
Since each entry $j\in\mathbb{I}_{6}$ occurs $12$ times in $\left[\begin{matrix}\mathbf{T}^{(1)}_{i}(1)\\ \mathbf{T}^{(2)}_{i}(1)\end{matrix}\right]$, $i=0,1,2$, which implies
$p_{X_{1}}(j)=\frac{12p_{0}}{72}+\frac{12p_{1}}{72}+\frac{12p_{2}}{72}=\frac{1}{6},$ (III.4.52)
and so
$H(X_{1})=\log 6=a+b.$ (III.4.53)
Similarly,
$H(X_{2})=H(X_{3})=\log 6=a+b.$ (III.4.54)
Each row of $\mathbf{T}(1,4)$ appears exactly once in $\mathbf{T}(1,4)$, which implies $H(X_{2},X_{3}|X_{1},X_{4})=0$. Hence,
$H(X_{1},X_{4})$ $=H(X_{1},X_{2},X_{3},X_{4})$ (III.4.55)
$=H(\underbrace{\frac{p_{0}}{72},\frac{p_{0}}{72},\dots,\frac{p_{0}}{72}}_{72},\dots,$ (III.4.56)
$\qquad\underbrace{\frac{p_{2}}{72},\frac{p_{2}}{72},\dots,\frac{p_{2}}{72}}_{72})$ (III.4.57)
$=H(p_{0},p_{1},p_{2})+\log{72}$ (III.4.58)
$=a-\log 2+\log 72$ (III.4.59)
$=3a+2b.$ (III.4.60)
By the same argument, we obtain
$H(X_{1234})=H(X_{1},X_{2},X_{3})=H(X_{1},X_{2},X_{4})$ (III.4.61)
$=H(X_{1},X_{3},X_{4})=H(X_{2},X_{3},X_{4})$ (III.4.62)
$=H(X_{1},X_{4})=H(X_{2},X_{4})=H(X_{3},X_{4})$ (III.4.63)
$=3a+2b.$ (III.4.64)
Each row of $\mathbf{T}(1,2)$ appears exactly twice in $\left[\begin{matrix}\mathbf{T}^{(1)}_{i}(1,2)\\ \mathbf{T}^{(2)}_{i}(1,2)\end{matrix}\right]$, $i=0,1,2$, which implies that
$p_{X_{1}X_{3}}(x_{1},x_{2})=\frac{2p_{0}}{72}+\frac{2p_{1}}{72}+\frac{2p_{2}}{72}=\frac{1}{36},$ (III.4.65)
and so
$H(X_{1},X_{2})=\log 36=2a+2b.$ (III.4.66)
By the same argument,
$H(X_{1},X_{2})=H(X_{1},X_{3})=H(X_{2},X_{3})=2a+2b.$ (III.4.67)
Then $(X_{i},i\in N_{4})$ characterizes $(a,b)$. The proof is accomplished. ∎
III-E Discussion
In this section, we characterized $10$ types of $2$-dimensional faces of $\Gamma_{4}$. For a face $F$ with one of its two extremes rays containing a rank-1 matroid, we can see that entropy region $F^{*}$ has a similar shape to the entropy region on $(U_{2,3},U^{12,3}_{1,2})$ characterized by Matúš in [10] or $(U_{2,3},U^{1,3}_{1,1})$ by Chen and Yeung in [11]. The characterization of such faces are generalized to faces of general $n$ and both extreme rays containing matroids in [32].
As each $\mathrm{MVOA}(P,v)$ with $M$ a rank 2 or greater integer polymatroid corresponds to a type of orthogonal Latin hypercubes, the characterization of a face with both extreme rays containing an integer polymatroid exceeding 1, in Subsection III-B-III-D, breeds a new combinatorial structure which can be considered as an intermediate form of the two types of orthogonal Latin hypercubes. Specifically, in Subsection III-C, three faces are characterized by the Latin square decompositions, which can be considered as three new types of Latin square substructures.